### v1

# Low-dimensional tori in Calogero–Moser–Sutherland systems

Abstract.
We give an explicit description of low-dimensional Liouville tori in Calogero–Moser–Sutherland systems.
1. Introduction
Recall that Calogero–Moser–Sutherland system [1, 6, 8] is a model of $n$ one-dimensional particles on a circle with trigonometric potential. The phase space of this model is $T^{*}H_{\rm reg}$ with coordinates $(p_{i},\gamma_{i})$, $p_{i}\in\mathbb{R}$, $\lvert\gamma_{i}\rvert=1$, $\sum_{i=1}^{n}p_{i}=0$, $\prod_{i=1}^{n}\gamma_{i}=1$.
The symplectic structure is standard with Poisson brackets:
(1) $\{p_{i},p_{j}\}=0,\quad\{p_{i},q_{j}\}=\delta_{ij},\quad\{q_{i},q_{j}\}=0.$
Here we used coordinates $q_{j},\gamma_{j}=e^{iq_{j}}$.
The Hamiltonian of the system is
(2) $H_{2}=\frac{1}{2}\sum_{i=1}^{n}p_{i}^{2}+\frac{1}{4}\sum_{i\neq j}^{n}\frac{c^{2}}{\sin^{2}(q_{i}-q_{j})}.$
This is an integrable system with higher integrals
(3) $H_{k}=\frac{1}{k}\mathrm{Tr}(x^{k}),$
where $x$ is given by formula (12). Function $H_{k}$ are invariant with respect to simultenious permutations of $p$ and $q$. Thus, effective phase space is an orbifold $T^{*}H_{reg}/S_{n}$, where $S_{n}$ acts on $\mathfrak{h}^{*}\times H_{reg}$ diagonally.
The main result of this paper is the following explicit description of all Liouville tori. Let $(x_{1},\dots,x_{n})$ be the eigenvalues of $x$ (given (12)). Then:
  • when $x_{i}-x_{i+1}>c$ for all $i$, the level surface of $\left\{H_{k}\right\}_{k=2}^{n}$ is $(n-1)$-dimensional and with well-known angle coordinates (see section 2.3).
  • when $x_{k_{i}}-x_{k_{i}+1}=c$ for some $1\leq k_{1}<\dots<k_{\ell}\leq n$, then the level surface of $\left\{H_{k}\right\}_{k=2}^{n}$ is $(n-\ell-1)$-dimensional (the generic case corresponds to $\ell=0$), with the angle variables described in section 3.
Generic $(n-1)$-dimensional tori are well known and completely described long time ago [7]. When $\ell=n-1$, the level surface is zero-dimensional, and is a fixed point of the multitime evolution (see also [LRS]).
$x_{2}-x_{3}$$x_{1}-x_{2}$cc2222-dim1111-dim1111-dim00-dim
Figure 1. Dimensions of Liouville tori over shifted principal Weyl chamber in the $x$-plane for $G=SU(3)$.
This paper was inspired and motivated by the work [5] on hybrid integrable systems and on the upcoming paper [2] on stratified integrable systems.
Similar stratification exists in many other models, for example, in the elliptic CM system and in integrable systems on moduli spaces of flat connections. We will address this in a separate publication. Superintegrable systems and, in particular, superintegrable systems of spin CMS type are studied in [2].
Acknowledgment
We are grateful to L. Feher for stimulating discussions and pointing out reference [3], to K. Jiang, Z. Chen, and H. Xiao for many discussions on stratified integrable systems.
The work of A. L. was supported by the Beijing Natural Science Foundation (IS24006). The work of N. R. was supported by the Collaboration Grant ”Categorical Symmetries” from the Simons Foundation, by the Changjiang fund, and by the project 075-15-2024-631 funded by the Ministry of Science and Higher Education of the RF.
2. Action-angle variables in the Calogero–Moser–Sutherland system
2.1. The Calogero–Moser–Sutherland system for $SU(n)$.
Here we review well known facts about the Calogero–Moser (CM) system for $SU(n)$. This section is a well-known extension of [4] from $T^{*}\mathfrak{g}$ to $T^{*}G$. We included it for clarity of exposition. Elements of $\mathfrak{su}(n)$ and traceless $n\times n$ hermitian matrices and elements of $SU(n)$ are unimodular $n\times n$ unitary matrices.
The phase space for the Calogero–Moser system for $G=SU(n)$(and of its spin analogs) can be realized as the symplectic reduction of $T^{*}G\simeq\mathfrak{g}\times G$ with respect to the Hamiltonian action of $G$ by conjugations. From now we identify $\mathfrak{g}\simeq\mathfrak{g}^{*}$ via the Killing form. For $\mathfrak{su}(n)$ this means that both $\mathfrak{su}(n)$ and $\mathfrak{su}(n)^{*}$ are represented by traceless $n\times n$ matrices. If $\mathcal{O}$ is a coadjoint orbit in $\mathfrak{su}(n)$ of rank one, the corresponding Hamiltonian reduction is
(4) $\mathcal{S}(\mathcal{O})=\{(x,g)\in\mathfrak{su}(n)\times SU(n)\mid x-g^{-1}xg\in\mathcal{O}\}/G,$
In general, it is a stratified symplectic space. When $\mathcal{O}$ is a rank $1$ orbit, it is isomorphic to $T^{*}H_{\mathrm{reg}}/S_{n},$ where $H_{\mathrm{reg}}=\{\mathrm{diag}(\gamma_{1},\ldots,\gamma_{n});\ \gamma_{i}\neq\gamma_{j}\}$ and is a symplectic orbifold. Since the Hamiltonians are also defined on a smooth manifold $T^{*}H_{reg}$, it is convenient to work there. We will denote it $\widetilde{\mathcal{S}(\mathcal{O})}$.
A rank $1$ orbit in $\mathfrak{su}(n)$ consists of traceless Hermitian matrices of rank $1$. They can be parametrized as
$\mu_{ij}=\overline{\varphi_{i}}\varphi_{j}-c\delta_{ij},\quad c=\frac{1}{n}\sum_{i=1}^{n}|\varphi_{i}|^{2}$
modulo transformations $\varphi_{j}\mapsto e^{i\theta}\varphi_{j}$. It is clear that the second Casimir function 111 The values of higher Casimir functions are the rank $1$ orbit are determinated by $c$ and (5) $\mathrm{Tr}(\mu^{k})=c^{k}\,(n-1)\left((n-1)^{k-1}+(-1)^{k}\right).$ is
$\mathrm{tr}(\mu^{2})=c^{2}n(n-1).$
The moment map brings the $SU(n)$-orbit through $(x,g)\in T^{*}G$ to
(6) $\mu=x-g^{-1}xg.$
From the construction of the phase space as a Hamiltonian reduction of $T^{*}SU(n)$ with respect to rank $1$ orbit the functions
(7) $H_{k}=\frac{1}{k}\mathrm{Tr}\big{(}x^{k}\big{)},\ \ k=2,\dots,n$
on $T^{*}G=\{(x,g)\in\mathfrak{su}(n)\times SU(n)$ Poisson commute. This follows immediately from the symplectic structure on $T^{*}G=\{(x,g)\in\mathfrak{su}(n)\times SU(n)$. As a consequence the Poisson commute on $\mathcal{S}(\mathcal{O})$. Functions $\{H_{k}\}_{k=2}^{n}$ are independent on $S(\mathcal{O})$ and because $\mathrm{dim}\,S(\mathcal{O})=2n-2$ they define a Liouville integrable system.
2.2. The diagonalization of $g$ and natural coordinates
Let us briefly recall the Hamiltonian reduction construction of Calogero-Moser system by Kazhdan, Kostant, and Sternberg [4], extended to Calogero-Moser-Sutherland. We can choose a representative of the $SU(n)$–orbit through $(x,g)$ in (4) such that $g$ is diagonal:
(8) $g=\mathrm{diag}(\gamma_{1},\ldots,\gamma_{n}).$
The equation (6) for diagonal $g$ becomes
(9) $x_{ij}(1-\gamma_{i}^{-1}\gamma_{j})=\overline{\varphi}_{i}\varphi_{j},\quad$ $i\neq j,$
(10) $0=|\varphi_{i}|^{2}-c,\quad\qquad\qquad$ $i=j.$
Thus $|\varphi_{i}|^{2}=c,\ c>0$.
Note that if $\gamma_{i}=\gamma_{j}$, the equation (9) implies $\overline{\varphi}_{i}\varphi_{j}=0$ but this contradicts to $\lvert\varphi_{i}\rvert^{2}=c$. Thus $\gamma_{i}\neq\gamma_{j}$.
When $\gamma_{i}\neq\gamma_{j}$ the stabilizer of (8) is the Cartan subgroup of $SU(n)$ of unitary diagonal unimodular matrices. By the action of this subgroup on $x$ we can bring $\varphi_{j}$ to the form $\varphi_{j}=\sqrt{c}\,e^{i\theta}$. This gives the description of the phase space as the quotient of $T^{*}H_{reg}$ (with coordinates $(x_{ii},\gamma_{i})$) with respect to the action of $S_{n}$ by permutation of pairs $(x_{ii},\gamma_{i})$.
For $j\neq k$ we have
(11) $x_{jk}=\frac{c}{1-\gamma_{j}^{-1}\gamma_{k}}.$
As it should be, we have $\bar{x}_{kj}=x_{jk}$.
Setting $p_{i}:=x_{ii}$ we have
(12) $x=\begin{pmatrix}p_{1}&\frac{c}{1-\gamma_{1}^{-1}\gamma_{2}}&\ldots&\frac{c}{1-\gamma_{1}^{-1}\gamma_{n}}\\ \frac{c}{1-\gamma_{2}^{-1}\gamma_{1}}&p_{2}&\ldots&\ldots\\ \ldots&\ldots&\ldots&\ldots\\ \frac{c}{1-\gamma_{n}^{-1}\gamma_{1}}&\ldots&\ldots&p_{n}\end{pmatrix}.$
The unimodularity implies $\sum_{i=1}^{n}p_{i}=0$. For the first two Hamiltonians we have:
(13) $H_{2}=\frac{1}{2}\mathrm{Tr}\big{(}x^{2}\big{)}=\frac{1}{2}\sum_{i=1}^{n}p_{i}^{2}+\frac{1}{2}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}\frac{c^{2}}{(1-\gamma_{j}^{-1}\gamma_{i})(1-\gamma_{i}^{-1}\gamma_{j})}=\frac{1}{2}\sum_{i=1}^{n}p_{i}^{2}+\frac{1}{4}\sum_{\begin{subarray}{c}i<j\end{subarray}}^{n}\frac{c^{2}}{\sin^{2}(\tfrac{q_{i}-q_{j}}{2})},$
(14) $H_{3}=\frac{1}{3}\mathrm{Tr}\big{(}x^{3}\big{)}=\frac{1}{3}\sum_{i=1}^{n}p_{i}^{3}+\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}\frac{c^{2}\,p_{i}}{(1-\gamma_{j}^{-1}\gamma_{i})(1-\gamma_{i}^{-1}\gamma_{j})}=\frac{1}{3}\sum_{i=1}^{n}p_{i}^{3}+\frac{1}{4}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}\frac{c^{2}\,p_{i}}{{\sin^{2}(\tfrac{q_{i}-q_{j}}{2})}}.$
Here $\gamma_{j}=\exp(iq_{j})$.
2.3. The diagonalization of $x$ and the action-angle variables
Geometrically, any integrable system is a projection from a $2n$-dimensional symplectic manifold to an $n$-dimensional space, such that the generic fiber is Lagrangian.
For Calogero–Moser systems we have the projection
$\pi\colon T^{*}H_{\mathrm{reg}}\to\mathfrak{h}^{*}/W,\quad(x,\gamma)\mapsto[x]\in\mathfrak{g}^{*}/G\simeq\mathfrak{h}^{*}/W.$
Here $x$ is given by (12) as the function of $p$ and $\gamma$, and for $\mathfrak{su}(n)$. $W=S_{n}$, and $[x]$ is a collection of eigenvalues of $x$. The quotient space $\mathfrak{h}^{*}/W$ can be modeled as:
$\mathfrak{h}^{*}/W\simeq\left\{x\in\mathbb{R}^{n}\ \middle|\ x_{1}+\dots+x_{n}=0,\ x_{1}\geq\dots\geq x_{n}\right\}.$
The main observation of this paper is that the image of this mapping is a shifted cone in $\mathbb{R}^{n-1}$. Preimages of strata of this cone stratify $T^{*}H_{\mathrm{reg}}$. For CMS model we describe the strata later in this section.
2.3.1. The generic stratum
Any $x\in\mathfrak{su}(n)^{*}$ is diagonalizable by the conjugation action of $SU(n)$. let us choose a representative $(x,g)$ in our phase space where $x$ is diagonal:
(15) $x=\mathrm{diag}\left(x_{1},\ldots,x_{n}\right)\in\mathfrak{h}^{*}\subset\mathfrak{su}(n)^{*},\qquad x_{1}+x_{2}+\ldots+x_{n}=0,\qquad x_{1}\geq x_{2}\geq\dots\geq x_{n}.$
The Hamiltonians, evaluated in terms of eigenvalues of $x$ are simply symmetric power functions:
$H_{k}=\sum_{j=1}^{n}x_{j}^{k}.$
Theorem 1.
Solutions to the momentum map equation
(16) $x-g^{-1}xg=\bar{\psi}\psi^{t}-c\cdot 1.$
can be described in the following way. Let’s $\lambda=g\bar{\psi}$, then the following holds
1) The following conditions are equivalen
  • $x_{i}-x_{i+1}\geq c$ for all $i=1,\ldots,n-1$,
  • $|\psi_{i}|^{2}\geq 0$ for all $i=2,\ldots,n$,
  • $|\lambda_{i}|^{2}\geq 0$ for all $i=1,\ldots,n-1$.
Moreover, $|\psi_{1}|^{2}>0$ and $|\lambda_{n}|^{2}>0$.
2) For any $n>k\geq 1$, the following occurs only at the same time
  • $x_{k}-x_{k+1}=c$,
  • $|\psi_{k+1}|^{2}=0$,
  • $|\lambda_{k}|^{2}=0$.
3) If $x_{i}-x_{i+1}>c$ for $i=1,\ldots,n-1$, then
(17) $|\psi_{i}|^{2}=c\prod_{j=1,j\neq i}^{n}\frac{x_{i}-x_{j}+c}{x_{i}-x_{j}},\qquad\text{and }\qquad|\lambda_{i}|^{2}=c\prod_{j=1,j\neq i}^{n}\frac{x_{j}-x_{i}+c}{x_{j}-x_{i}}.$
4) For $x_{i}-x_{i+1}>c$ with $i=1,\ldots,n-1$ matrix $g$ has form
(18) $g_{ij}=\frac{\lambda_{i}\,\psi_{j}}{x_{j}-x_{i}+c},$
Proof.
1) The equation for the moment map can be written as
(19) $g^{-1}xg=x+c\cdot 1-\bar{\psi}\psi^{t},$
Following lemma 3.4 from [3], let’s consider the equality of characteristic polynomials of both parts of the equation (31)
(20) $\prod_{j=1}^{n}(x_{j}-z)=\mathrm{det}(x+c\cdot 1-\bar{\psi}\psi^{t}-z)=\prod_{j=1}^{n}(x_{j}+c-z)-\sum_{j=1}^{n}|\psi_{j}|^{2}\prod_{k\neq j}(x_{k}+c-z).$
Here we take into account that for diagonal matrix $A$
(21) $\mathrm{det}\left(A+uv^{t}\right)=\mathrm{det}\left(A\right)+v^{t}A^{adj}u,$
where $A^{adj}$ is the adjugate matrix to $A$, which defined by $A^{adj}A=\mathrm{det}A\cdot 1$. This is a well-known fact which follows from Lemma 1 in the appendix A.
Equation (20) at the point $z=x_{i}+c$ is
(22) $\prod_{j=1}^{n}(x_{i}-x_{j}+c)=|\psi_{i}|^{2}\prod_{j\neq i}(x_{i}-x_{j}).$
It follows from the equation (22) for $i=1$ that $|\psi_{1}|$ is strictly positive
(23) $|\psi_{1}|^{2}=c\prod_{j>1}\frac{x_{1}-x_{j}+c}{x_{1}-x_{j}}>0.$
In equation (22) for $i=2$
(24) $c\ \frac{(x_{2}-x_{1}+c)}{x_{2}-x_{1}}\prod_{j>2}\frac{(x_{2}-x_{j}+c)}{(x_{2}-x_{j})}=|\psi_{2}|^{2}$
the product over $j>2$ is strictly positive. So we have
(25) $|\psi_{2}|^{2}\geq 0\qquad\Leftrightarrow\qquad x_{1}\geq x_{2}+c,$
where equality $x_{1}=x_{2}+c$ is achieved if and only if $|\psi_{2}|^{2}=0$.
Continuing this reasoning, we get the statement of the proposition. So, we have
(26) $|\psi_{k}|^{2}\geq 0\qquad\Leftrightarrow\qquad x_{k-1}\geq x_{k}+c,$
where equality $x_{k-1}=x_{k}+c$ is achieved if $|\psi_{k}|^{2}=0$.
The equation for the moment map can be rewritten for $\lambda$ as
(27) $gxg^{-1}=x-c\cdot 1+\lambda\bar{\lambda}^{t},$
Equality of characteristic polynomials gives
(28) $\prod_{j=1}^{n}(x_{j}-z)=\mathrm{det}(x-c\cdot 1+\lambda\bar{\lambda}^{t}-z)=\prod_{j=1}^{n}(x_{j}-c-z)+\sum_{j=1}^{n}|\lambda_{j}|^{2}\prod_{k\neq j}(x_{k}-c-z).$
Equation (28) at the point $z=x_{i}-c$ is
(29) $\prod_{j=1}^{n}(x_{j}-x_{i}+c)=|\lambda_{i}|^{2}\prod_{j\neq i}(x_{j}-x_{i}).$
Using a similar argument as for equation (22), we get $|\lambda_{n}|^{2}>0$ and
(30) $|\lambda_{k-1}|^{2}\geq 0\qquad\Leftrightarrow\qquad x_{k-1}\geq x_{k}+c,$
where equality $x_{k-1}=x_{k}+c$ is achieved if $|\lambda_{k-1}|^{2}=0$.
2) This part of the statement directly follows from (22) and (29).
3) After the multiplication of the momentum map (31) by $g$ the equation for the moment map becomes
(31) $gx+cg-xg=\lambda\psi^{t},$
$(i,j)$ whose element has the form
(32) $(x_{j}-x_{i}+c)g_{ij}=\lambda_{i}\psi_{j},$
which fix the form of the thirt part of the statement
which implies (18).
Remark 1.
If $x_{i}>x_{i+1}+c$, the stabilizer of $x$ is the Cartan subgroup in $SU(n)$, i.e., unitary diagonal matrices. Acting by this stabilizer on $\psi$, we can bring it to the form $\psi_{i}=|\psi_{i}|$ for all $i$. After this, we have:
(33) $g=\alpha X,\qquad X_{ij}=\frac{|\lambda_{i}||\psi_{j}|}{x_{j}-x_{i}+c},\qquad\alpha_{ij}=\alpha_{i}\delta_{ij},\qquad|\alpha_{i}|=1,$
where $X$ is orthogonal matrix, i.e. $X^{t}=X^{-1}$. with $\mathrm{det}\,X=1$ according to
The Cauchy determinant formula implies that $\mathrm{det}\,X=1$, and therefore $\alpha_{1}\cdot\ldots\cdot\alpha_{n}=1$.
Remark 2.
We can multiply $g$ by a central element $\omega\in Z(SU(n))$:
(34) $\omega=\exp\left(\frac{2\pi i}{n}k\right)\cdot 1_{n},\qquad\text{ where }k=0,\ldots,n-1.$
If $\gamma_{j}$ are eigenvalues of the matrix $g$, such multiplication changes $\gamma_{j}\mapsto\omega\gamma_{j}$. Eigenvalues of $g$ are exponents of coordinates of Calogero–Moser–Sutherland model $\gamma_{j}=\exp(iq_{j})$. Thus the multiplication of $g$ by $\omega$ shifts coordinates as
[…]
Thus, we described coordinates on the generic (maximal) stratum: real coordinates $\{x_{i}\}$ and $\{\mathrm{arg}\,\alpha_{i}\}$ with satisfying constrains $x_{i}>x_{i+1}+c,\ \sum_{i=1}^{n}x_{i}=0$ and $\prod_{i=1}^{n}\alpha_{i}=1$, $|\alpha_{i}|=1$.
Now let us describe all strata and let us show that $(x,\alpha)$ are the action-angle variables.
Define $y_{i}=x_{i-1}-x_{i}$ for $j=2,\ldots,n$, such that $y_{i}\geq c$. Then
(36) $x_{i}=\frac{1}{n}\sum_{j=2}^{n}(n+1-j)y_{j}-\sum_{j=2}^{i}y_{j}.$
Define also $\theta_{j}=\tfrac{1}{i}\sum_{k=1}^{j-1}\log\alpha_{k}$ for $j=2,\ldots,n$. Then,
(37) $\alpha_{k}=e^{i(\theta_{k+1}-\theta_{k})},\qquad k=1,\ldots n.$
Here we assume that $\theta_{1}=\theta_{n+1}=0$.
Thus, we get coordinates on the generic (maximal) stratum: real coordinates $(y_{2}\ldots,y_{n})$ satisfying constrains $y_{i}>c$ and $(\theta_{2}\ldots,\theta_{n})$. In section 4 we show that $(y,\theta)$ are the action-angle variables.
3. Low–dimensional tori
The moment map equation (6) when $x$ is diagonal becomes
(38) $(x_{j}-x_{i}+c)g_{ij}=\lambda_{i}\psi_{j},$
where $\lambda=g\bar{\psi}$.
Proposition 1.
If for some $k$ we have $\psi_{k}=0$, then in the $k$-th column and $(k-1)$-th row of the martrix $g$ only nonzero element is $g_{k-1,k}\neq 0$ and $|g_{k-1,k}|=1$.
Proof.
Consider equation (38) for $j=k$
(39) $(x_{k}-x_{i}+c)g_{ik}=0.$
If for all $i$ elements $g_{ik}=0$ then $g$ not invertible. Theorem 1 implies that $\psi_{k}=0$ if and only if $x_{k}-x_{k-1}+c=0$. For $j\neq k-1$ $x_{j}-x_{k}+c\neq 0$ and therefore $g_{jk}=0$. This means that $k$-th column has only one non-zero element $g_{k-1,k}$.
Unitarity condition $\left(g^{{\dagger}}g\right)_{ik}=\sum_{j=1}^{n}\bar{g}_{ji}g_{jk}=\bar{g}_{k-1,i}g_{k-1,k}=\delta_{ik}$ implies that in the $(k-1)$-th row there is only one non zero element $g_{k-1,k}$, and $|g_{k-1,k}|=1$. ∎
Assume that $\psi_{j}\neq 0$ only for $j=1,j_{1},\dots,j_{s}$ where $j_{1}<j_{2}<\dots<j_{s}$. For convenience, define $j_{0}=1$.
Define the matrix $\widehat{g}$ as $P\widehat{g}=g$, where $P$ is the matrix of cyclic permutation with sign
(40) $P=\begin{pmatrix}0&-1&0&\ldots&0\\ 0&0&-1&\ldots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\ldots&-1\\ 1&0&0&\ldots&0\\ \end{pmatrix}.$
Proposition 1 implies
$\widehat{g}_{ij}=\delta_{ij}\alpha_{i},\quad|\alpha_{i}|=1$
if neither $i$ no $j$ are not in $\{j_{a}\}_{a=0}^{s}$.
The moment map equations give the following identity
$x-\widehat{g}^{-1}x^{P}\widehat{g}=\overline{\psi}\psi^{t}-c$
where $x_{j}^{P}=x_{j-1}$ with $x_{0}=x_{n}$, and equations
(41) $\left(x_{j_{b}}-x^{P}_{j_{a}}+c\right)\widehat{g}_{j_{a}j_{b}}=\sum_{d=0}^{s}\widehat{g}_{j_{a}j_{d}}\,\overline{\psi}_{j_{d}}\,\psi_{j_{b}}$
Define
$\widehat{\lambda}_{j_{a}}=\sum_{d=0}^{s}\widehat{g}_{j_{a}j_{d}}\,\overline{\psi}_{j_{d}}.$
Theorem 2.
The following holds:
  • $\widehat{\lambda}_{i}\neq 0,\lambda_{i}$, for $i\neq j_{a}$, $a=0,\dots,s$.
  • (42) $|\psi_{j_{a}}|^{2}=\frac{\prod_{d=0}^{s}(x_{j_{a}}-x^{P}_{j_{d}}+c)}{\prod^{s}_{d\neq a}(x_{j_{a}}-x_{j_{d}})},$
  • (43) $|\widehat{\lambda}_{j_{a}}|^{2}=\frac{\prod_{d=0}^{s}(x_{j_{d}}-x^{P}_{j_{a}}+c)}{\prod^{s}_{d\neq a}(x^{P}_{j_{d}}-x^{P}_{j_{a}})}$
  • By the action of the Cartan part of the stabilizer of $x$, we can bring $\widehat{g}$ to the form
    (44) $\widehat{g}_{j_{a}j_{b}}=\alpha_{j_{a+1}}\frac{|\widehat{\lambda}_{j_{a}}|\,|\psi_{j_{b}}|}{x_{j_{b}}-x^{P}_{j_{a}}+c}$
Proof.
Write the moment map equation as
$\hat{g}^{-1}x^{P}\hat{g}=x-\overline{\psi}\psi^{t}+c.$
For the characteristic polynomial of the right side, using lemma 1 from the appendix A we obtain
$\prod_{d=0}^{s}(x_{j_{d}}^{P}-z)=\prod_{d=0}^{s}(x_{j_{d}}+c-z)-\sum_{b=0}^{s}|\psi_{j_{b}}|^{2}\prod_{\begin{subarray}{c}d=0\\ d\neq b\end{subarray}}^{s}(x_{j_{d}}+c-z).$
Taking $z=x_{j_{a}}+c$ we obtain
$\prod_{d=0}^{s}(x_{j_{d}}^{P}-x_{j_{a}}-c)=-|\psi_{j_{a}}|^{2}\prod_{\begin{subarray}{c}d=0\\ d\neq a\end{subarray}}^{s}(x_{j_{d}}-x_{j_{a}})$
which proves the formula (42) for $|\psi_{j_{a}}|^{2}$.
Similarly, writing the moment map equation as
$\hat{g}\,x\,\hat{g}^{-1}=x^{P}+\widehat{\lambda}\,\widehat{\lambda}^{*}-c$
and using lemma 1 gives the proof of the formula (43) for $|\widehat{\lambda}_{j_{a}}|^{2}$.
Finally, the moment map equations (41) give
$\widehat{g}_{j_{a}j_{b}}=\frac{\widehat{\lambda}_{j_{a}}\,\psi_{j_{b}}}{x_{j_{b}}-x_{j_{a}}^{P}+c}.$
By the action of the stabilizer of $x$, which is the Cartan subgroup $SU(n)$, we can make $\psi_{j_{b}}$ real. This gives (44). ∎
Note that to make $\psi_{j_{a}}$ real we need only a subgroup of the stabilizer.
Remark 3.
We can write $g$ as the product
(45) $g=\alpha\,P\,Y,$
where $P$ is the cyclic permutation matrix (40).
Let $\beta$ be an element of the stabilizer $x$, i.e. a diagonal unitary matrix.
We use transformations
(46) $g\to\beta g\beta^{-1}$
with $\beta_{i}=1$, when $i\neq j_{a}$, to make $\psi_{j_{a}}$ real.
Let $\beta^{\prime}\in H$ be such that $\beta^{\prime}_{i}=1$, when $i=j_{a}$. We have
$\beta^{\prime}g\beta^{\prime-1}=\beta^{\prime}\alpha PY\beta^{\prime-1}=\beta^{\prime}\alpha\beta^{\prime-1}_{P}PY.$
We can use these transformations to bring $\alpha$ to the form $\alpha_{i}=1$, $i\neq j_{a}$.
Thus, we can fix the section of the action of the stabilizer group $H\subset SU(n)$:
(47) $g=\alpha\,P\,Y,$
where
$Y_{ij}$ $=\delta_{ij},\quad i,j\notin\{j_{0},\ldots,j_{s}\}$
(48) $Y_{i,j_{a}}=Y_{j_{a},i}$ $=0,\quad i\notin\{j_{0},\ldots,j_{s}\}$
(49) $Y_{j_{a}j_{b}}$ $=(-1)^{1-\delta_{a,0}}\frac{|\widehat{\lambda}_{j_{a}}||\psi_{j_{b}}|}{x_{j_{b}}-x_{j_{a}}^{P}+c},$
(50) $\alpha_{ij}=\alpha_{i}\delta_{ij},\qquad|\alpha_{i}|=1,\qquad\text{and}\quad\alpha_{i}=1\ \text{ for }\ i\neq j_{a},$
and $P$ is defined (40).
Note that with this choice, we have
(51) $\mathrm{det}\,\alpha=\mathrm{det}\,P=\mathrm{det}\,Y=1$
and real matrix $Y$ is othogonal
$YY^{t}=I.$
Define $y_{j_{a}}=x_{j_{a}}-x_{j_{a}-1},$ then due to (36)
(52) $x_{i}=\frac{n-1}{2}c-(i-1)c+\sum_{a=1}^{s}\left(\frac{n+1-j_{a}}{n}-\delta_{i\geq j_{a}}\right)(y_{j_{a}}-c).$
Thus, we described coordinates $(y_{j_{a}},\alpha_{j_{a}})$ on $s$-dimention strata: $y_{j_{a}}>c$ and $\{\alpha_{j_{a}}\}_{a=0}^{s}$ with constrain $\prod_{a=0}^{s}\alpha_{j_{a}}=1$.
Example 1.
(The fixed point). In the case when all $\psi_{i}=0$ for $i\geq 2$, we have
  • $|\psi_{1}|^{2}=|\hat{\lambda}_{1}|^{2}=nc$.
  • $x_{i}=\frac{n-1}{2}c-(i-1)c$ for $1=2,\ldots,n$.
(53) $g=P.$
In terms of original CMS coordinates we have $q_{j}=\frac{2\pi j}{n}$. The action of the center of $SU(n)$, $g\mapsto\omega^{k}g$, $\omega^{n}=1$, acts by the cyclic shift $q_{j}\mapsto q_{j}+\frac{2\pi k}{n}\mod(2\pi)$.
Example 2.
(One-domensional tori). Fix $2\leq j\leq n$ and assume that $\psi_{i}=0$ when $i\neq 1,j$.
  • $|\psi_{1}|^{2}+|\psi_{j}|^{2}=|\hat{\lambda}_{1}|^{2}+|\hat{\lambda}_{j}|^{2}=nc$.
  • For $y=x_{j-1}-x_{j}$, $y>c$; $x_{i-1}-x_{i}=c$ for $i=2,\ldots,n$ and $i\neq j$, and with $\sum_{i=1}^{n}x_{i}=0$:
    (54) $x_{i}=\frac{n-1}{2}c-(i-1)c+\left(\frac{n+1-j}{n}-\delta_{i\geq j}\right)(y-c),$
    where $\delta_{i\geq j}$ is a step function.
  • $g=\alpha PY$, where $\alpha=\mathrm{diag}\left(\alpha_{1},1,\ldots,1,\alpha_{j},1,\ldots,1\right)$.
(55) $|\psi_{1}|^{2}$ $=(j-1)c\,\frac{y+(n-1)c}{y+(j-2)c},\qquad|\psi_{j}|^{2}=(n-j+1)c\,\frac{y-c}{y+(j-2)c},$
$|\hat{\lambda}_{1}|^{2}$ $=(n-j+1)c\frac{y+(n-1)c}{y+(n-j)c},\qquad|\hat{\lambda}_{j}|^{2}=(j-1)c\ \frac{y-c}{y+(n-j)c}$
and
(56) $Y=\begin{pmatrix}r&0&\ldots&q&\ldots&0\\ 0&1&\ldots&0&\ldots&0\\ \vdots&\vdots&\ddots&\vdots&\ddots&\vdots\\ -q&0&\ldots&r&\ldots&0\\ \vdots&\vdots&\ddots&\vdots&\ddots&\vdots\\ 0&0&\ldots&0&\ldots&1\\ \end{pmatrix},$
where
(57) $r^{2}=\frac{(j-1)c\,(n-j+1)c}{(y+(j-2)c)(y+(n-j)c)},\qquad q^{2}=\frac{(y-c)(y+(n-1)c)}{(y+(j-2)c)(y+(n-j)c)}.$
Note that $r^{2}+q^{2}=1$. It is clear that $\mathrm{det}(\,Y)=1$, therefore $\mathrm{det}\,\alpha=\alpha_{1}\,\alpha_{j}=1$.
Remark 4.
Theorem 2 can be obtained from theorem 1 directly using the appropriate limit. In particular, if $x_{i-1}\to x_{i}+c$ for $i\neq j_{a}$, $a=0,\dots,s$ ($j_{0}=1$)
(58) $|\psi_{i}|^{2}$ $\to 0,\qquad\text{for }i\neq j_{a},$
(59) $|\psi_{i}|^{2}$ $\to|\psi_{j_{a}}|^{2},\qquad\text{for }i=j_{a},$
where in the LHS $|\psi_{i}|^{2}$ from (17) and in the RHS $|\psi_{j_{a}}|^{2}$ from (42), and
(60) $|\lambda_{i}|^{2}$ $\to 0,\qquad\text{for }i\neq j_{a}-1,$
(61) $|\lambda_{i}|^{2}$ $\to|\widehat{\lambda}_{j_{a}}|^{2},\qquad\text{for }i=j_{a}-1\ \mathrm{mod}\ n.$
4. Action-angle variables
The cotangent bundle $T^{*}G$ carries the natural symplectic form
(62) $\Omega=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}g\,g^{-1}\right).$
Here we used the trivialization $T^{*}G\simeq\mathfrak{g}^{*}\times G$ by right translation. The identification $\mathfrak{g}^{*}\simeq\mathfrak{g}$ is given by the Killing form. We also assume that $G$ is a compact matrix group.
This symplectic form is $G$-invariant with respect to the lift of the conjugation action to $T^{*}G$. After the trivialization and identification $\mathfrak{g}^{*}\simeq\mathfrak{g}$, it acts as
$h\colon\quad x\mapsto hxh^{-1},\quad g\mapsto hgh^{-1}.$
The symplectic form (62) restricted to $\mu^{-1}(\mathcal{O})$ gives a presymplectic form, which remains $G$-invariant, and becomes a symplectic form $\Omega_{\mathcal{O}}$ on $S(\mathcal{O})=\mu^{-1}(\mathcal{O})/G$. We can model $S(\mathcal{O})$ as the space of pairs $(x,g)$, up to the action $H\subset SU(n)$, where $x=(x_{1}\geq x_{2}\geq\ldots\geq x_{n})$, $\sum_{i=1}^{n}x_{i}=0$, is diagonalized. On the big stratum $g=\alpha X$ (33), on others strata $g=\alpha PY$, see remark 3.
Now, let us describe the restriction $\Omega_{\mathcal{O}}$ of the symplectic form (62) in terms of $(y,\theta)$ coordinates.
Proposition 2.
On the generic stratum
(63) $\Omega_{\mathcal{O}}=-\sum_{j=1}^{n}\mathrm{d}x_{j}\wedge\mathrm{d}\eta_{j}.$
where $x_{i}$ are eigenvalues of $x$ and $\alpha_{j}=\mathrm{exp}(i\eta_{j})$.
Proof.
Using the representation (33) for $\Omega$ form we get
(64) $\Omega_{\mathcal{O}}=\left.i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}g\,g^{-1}\right)\right|_{\mu^{-1}(\mathcal{O})}=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}(\alpha X)\,X^{-1}\alpha^{-1}\right)=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}X\,X^{-1}\right)+i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}\alpha\,\alpha^{-1}\right).$
The first term vanishes:
(65) $\mathrm{Tr}\left(x\,\mathrm{d}X\,X^{-1}\right)=\sum_{i,j=1}^{n}x_{i}\,\mathrm{d}X_{ij}\cdot(X^{t})_{ji}=\frac{1}{2}\sum_{i,j=1}^{n}x_{i}\,\mathrm{d}(X_{ij})^{2}=\frac{1}{2}\sum_{i=1}^{n}x_{i}\,\mathrm{d}\sum_{j=1}^{n}(X_{ij})^{2}=0,$
where we take into account the orthogonality of $X$, which implies $\sum_{j=1}^{n}(X_{ij})^{2}=1$.
Thus
(66) $\Omega_{\mathcal{O}}=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}\alpha\,\alpha^{-1}\right)=i\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\alpha^{-1}_{i}\mathrm{d}\alpha_{i}=-\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}\eta_{i}.$
Taking into account constraints $\sum_{j=1}^{n}\mathrm{d}x_{j}=0$ and $\sum_{j=1}^{n}\mathrm{d}\eta_{j}=0$ we can write $\Omega_{\mathcal{O}}$ as
(67) $\Omega_{\mathcal{O}}=-\sum_{j=2}^{n}\mathrm{d}y_{j}\wedge\mathrm{d}\theta_{j}.$
where $y_{j}=x_{j-1}-x_{j}$ and $\theta_{j}=\sum_{i=1}^{j-1}\eta_{i}$. Thus, on the generic stratum $(y_{j},\theta_{j})$ are Darboux action-angle coordinates.
Proposition 3.
On a $2s$-dimensional stratum of $S(\mathcal{O})$ the symplectic form $\Omega_{\mathcal{O}}$ becomes
(68) $\Omega_{\mathcal{O}}=-\sum_{a=1}^{s}\mathrm{d}y_{j_{a}}\wedge\mathrm{d}\theta_{j_{a}}.$
where $y_{j_{a}}=x_{j_{a}-1}-x_{j_{a}}$ and $\theta_{j_{a}}=\sum_{b=a}^{s}\eta_{j_{b}}$, with $\alpha_{j_{b}}=e^{i\eta_{j_{b}}}$.
Proof.
Using representation (47) for $\Omega$ form we get
(69) $\Omega_{\mathcal{O}}=\left.i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}g\,g^{-1}\right)\right|_{\mu^{-1}(\mathcal{O})}=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}(\alpha PY)\,Y^{-1}P^{-1}\alpha^{-1}\right)=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}Y\,Y^{-1}\right)+i\,\mathrm{d}\mathrm{Tr}\left(x^{P}\,\mathrm{d}\alpha\,\alpha^{-1}\right),$
where $x^{P}=P^{-1}xP$.
Define $\alpha_{j_{a}}=e^{i\eta_{j_{a}}}$ for $a=0,\ldots,s$. Constrains $\prod_{a=0}^{s}\alpha_{j_{a}}=1$ implies.
(70) $\eta_{j_{0}}\equiv\eta_{1}=-\sum_{a=1}^{s}\eta_{j_{a}}.$
Expression $\mathrm{Tr}\left(x^{P}\,\mathrm{d}Y\,Y^{-1}\right)=0$ for the same reasons as in equation (65) and $\Omega_{\mathcal{O}}$ becomes
(71) $\Omega_{\mathcal{O}}=i\sum_{a=0}^{s}\mathrm{d}x^{P}_{j_{a}}\wedge\alpha_{j_{a}}^{-1}\mathrm{d}\alpha_{j_{a}}=\\ -\sum_{a=1}^{s}\frac{1-j_{a}}{n}\mathrm{d}y_{j_{a}}\wedge\mathrm{d}\eta_{1}-\sum_{a=1}^{s}\sum_{b=1}^{s}\left(\frac{n+1-j_{b}}{n}-\delta_{j_{b}-1\geq j_{a}}\right)\mathrm{d}y_{j_{b}}\wedge\mathrm{d}\eta_{j_{a}}=\\ -\sum_{a=1}^{s}\sum_{b=1}^{s}\left(1-\delta_{j_{b}-1\geq j_{a}}\right)\mathrm{d}y_{j_{b}}\wedge\mathrm{d}\eta_{j_{a}}=-\sum_{b=1}^{s}\mathrm{d}y_{j_{b}}\wedge\sum_{a=b}^{s}\mathrm{d}\eta_{j_{a}}.$
Here we used (70). In term of $\theta_{j_{a}}=\sum_{a=b}^{s}\eta_{j_{b}}$ this gives (68). ∎
Remark 5.
Note that $\theta$’s defined in proposition 3 differ slightly from $\theta$’s defined in proposition 2. This difference arises due to the introduction of a shift matrix $P$ (40).
The coordinates $(y_{j_{a}},\theta_{j_{a}})$ are Darboux action-angle coordinates on this $s$-dimensional stratum.
5. The multitime dynamics of Calogero–Moser–Sutherland system
5.1. The multi-time dynamics on $T^{*}G$.
Let $v_{k}$ be the Hamiltonian vector field on $T^{*}G\simeq\mathfrak{g}^{*}\times G$ generated by $H_{k}=\frac{1}{k}\operatorname{tr}(x^{k})$, i.e. $i_{v_{k}}\Omega=-dH_{k}$, where $\Omega$ is the symplectic form (62) on $T^{*}G$. We have
(72) $v_{k}=-i\sum_{i,j=1}^{n}\left(x^{k-1}g\right)_{ij}\frac{\partial}{\partial g_{ij}},$
Because Hamiltonians $H_{k}$ Poisson commute on $T^{*}G$, vector fields $v_{k}$ also commute. Their multi-time flows are solutions to the system of differential equations:
(73) $\frac{\mathrm{d}x({\bf t})}{\mathrm{d}t_{k}}$ $=0,$
(74) $\frac{\mathrm{d}g({\bf t})}{\mathrm{d}t_{k}}$ $=ix({\bf t})^{k-1}g({\bf t}).$
The initial value problem $x(0)=x,\,g(0)=g$ for these equations can be solved explicitly:
(75) $x({\bf t})=x,\qquad g({\bf t})=e^{i\sum_{k=1}^{n}t_{k}\,x^{k-1}}g.$
This dynamics preserves Hamiltonians $H_{k}$ and the moment map $\mu(x({\bf t}),g({\bf t}))=\mu(x,g)=x-g^{-1}xg$.
5.2. The multi-time dynamics on the generic stratum.
In coordinates $(y,\theta)$ the symplectic form on $S(\mathcal{O})$ obtained by the Hamiltonian reduction from (62) is
(76) $\Omega_{\mathcal{O}}=-\sum_{j=2}^{n}\mathrm{d}y_{j}\wedge\mathrm{d}\theta_{j},$
which means Poisson brackets between coordinate function $\{y_{i}\}$ and $\{\theta_{i}\}$ are
$\{y_{i},y_{j}\}=0,\qquad\{\theta_{i},\theta_{j}\}=0,\qquad\{\theta_{i},y_{j}\}=\delta_{ij}.$
Any $G$-invariant function $H$ on $T^{*}G$ defines a $G$-invariant function on $\mu^{-1}(\mathcal{O})\subset T^{G}$ and a function $S(\mathcal{O})=\mu^{-1}(\mathcal{O})/G$ which we denote by same letter $H$. Hamiltonian vector fields $v_{k}$ restrict to $S(\mathcal{O})$ and have the form
(77) $v_{k}=\sum_{i,j}x_{i}^{k-1}\frac{\partial x_{i}}{\partial y_{j}}\frac{\partial}{\partial\theta_{j}},\qquad\text{ for }k=2,\ldots,n.$
Here we can express $x_{i}$ in terms of $y_{i}$ according to (36):
(78) $v_{k}=\sum_{j=2}^{n}u^{k}_{j}\frac{\partial}{\partial\theta_{j}},$
where
(79) $u^{k}_{j}=\sum_{i=1}^{n}\left(\frac{1}{n}\sum_{j=2}^{n}(n+1-j)y_{j}-\sum_{j=2}^{i}y_{j}\right)^{k-1}\left(\frac{n+1-j}{n}-\delta_{i\geq j}\right).$
Flow lines of these vector fields are linear in coordinates $\theta_{l}$. The multitime flow line with $y_{j}(0)=y_{j}$, $\theta_{j}(0)=\theta_{j}$ is
(80) $y_{j}({\bf t})=y_{j},\qquad\theta_{j}({\bf t})=\theta_{j}-\sum_{k=1}^{n-1}u^{k}_{j}\,t_{k},\qquad\text{ for }k=2,\ldots,n.$
From this we have:
(81) $x({\bf t})=x,\qquad g({\bf t})=\alpha(\mathbf{t})Y,$
where $\alpha_{k}(\mathbf{t})=\exp\left(i(\theta_{k+1}(\mathbf{t})-\theta_{k}(\mathbf{t}))\right)$ with $\theta_{1}(\mathbf{t})=\theta_{n+1}(\mathbf{t})=0$ and $Y$ is defined in (33).
The time dependence of coordinates $q_{j}$ is determined by $g(\mathbf{t})$ since $\{e^{iq_{j}(\mathbf{t})}\}$ are eigenvalues of $g(\mathbf{t})$.
5.3. The multi-time dynamics on low-dimensional strata.
When $\psi_{i}=0$ except for $i=j_{0},j_{1},\ldots,j_{s}$ (here $j_{0}=1$), the matrix $g$ can be written as
(82) $g=\alpha PY.$
In coordinates $(y,\theta)$ the symplectic form on the correspond stratum of $S(\mathcal{O})$, see remark 3, is
(83) $\Omega=-\sum_{a=1}^{s}\mathrm{d}y_{j_{a}}\wedge\mathrm{d}\theta_{j_{a}},$
which implies
$\{y_{j_{a}},y_{j_{b}}\}=0,\qquad\{\theta_{j_{a}},\theta_{j_{b}}\}=0,\qquad\{\theta_{j_{a}},y_{j_{b}}\}=\delta_{ab}.$
Hamiltonian vector fields $v_{k}$ restrict to $S(\mathcal{O})$ and have the form
(84) $v_{k}=\sum_{i=1}^{n}\sum_{a=1}^{s}x_{i}^{k-1}\frac{\partial x_{i}}{\partial y_{j_{a}}}\frac{\partial}{\partial\theta_{j_{a}}},\qquad\text{ for }k=2,\ldots,n.$
Here we can choose $s$ independent vector fields. In particular, we can choose the first $s$ vectors.
Here we can express $x_{i}$ in terms of $y_{i}$ according to (36):
(85) $v_{k}=\sum_{a=2}^{s}u^{k}_{a}\frac{\partial}{\partial\theta_{j_{a}}},$
where
(86) $u^{k}_{a}=\sum_{i=1}^{n}\left(\frac{n-1}{2}c-(i-1)c+\sum_{a=1}^{s}\left(\frac{n+1-j_{a}}{n}-\delta_{i\geq j_{a}}\right)(y_{j_{a}}-c)\right)^{k-1}\left(\frac{n+1-j_{a}}{n}-\delta_{i\geq j_{a}}\right).$
The solution to the initial value problem $y_{j_{a}}(0)=y_{j_{a}},\,\theta_{j_{a}}(0)=\theta_{j_{a}}$ for these equations is
(87) $y_{j_{a}}({\bf t})=y_{j_{a}},\qquad\theta_{j_{a}}({\bf t})=\theta_{j_{a}}-\sum_{k=1}^{s}u^{k}_{a}\,t_{k},\qquad\text{ for }a=1,\ldots,s.$
So,
(88) $x({\bf t})=x,\qquad g({\bf t})=\alpha(\mathbf{t})P\,Y,$
where $\alpha_{j_{a}}(\mathbf{t})=\exp\left(i(\theta_{j_{a}}(\mathbf{t})-\theta_{j_{a+1}}(\mathbf{t}))\right)$ with $\theta_{1}(\mathbf{t})=\theta_{j_{s+1}}(\mathbf{t})=0$, $Y$ is defined in (48) and $\alpha$ is defined in (50).
The time dependence of coordinates $q_{j}$ is determined by $g(\mathbf{t})$ since $\{e^{iq_{j}(\mathbf{t})}\}$ are eigenvalues of $g(\mathbf{t})$.
Example 3.
Zero-dimensional stratum.
On this stratum, particles are in equilibrium. Their position are terminated by eigenvalues of (53)
(89) $\mathrm{det}\left(\lambda-g\right)=\lambda^{n}+(-1)^{n}=0.$
Example 4.
One-dimensional strata.
There are $(n-1)$ such strata, labeled by $j=2,\ldots,n$, with coordinates $(y_{j},\theta_{j})$ defined in example 2.
Positions of particles are eigenvalues of $g(t)$ and thus are roots of the characteristic equation
(90) $\mathrm{det}\left(\lambda-g\right)=\lambda^{n}+(-1)^{j-1}q\beta^{-1}\lambda^{n-j+1}+(-1)^{n-j+1}q\beta\lambda^{j-1}+(-1)^{n}=0,$
where $\beta=\beta(t)=\alpha_{1}^{-1}=\alpha_{j}=\exp\left(i\theta_{j}(t)\right)=\exp(i\,(\theta_{j}-u\,t))$ and $10<q<$ 1 defined in example 2, and
(91) $u=\sum_{i=1}^{n}\left(\frac{n-1}{2}c-(i-1)c+\left(\frac{n+1-j}{n}-\delta_{i\geq j}\right)(y_{j}-c)\right)\left(\frac{n+1-j}{n}-\delta_{i\geq j}\right)=\\ (n-j+1)(j-1)\left(\frac{c}{2}+\frac{y-c}{n}\right).$
Appendix A Lemma about determinant of sum of matrix plus rank $1$ matrix
Lemma 1.
Let $D$ be a diagonal matrix $D_{ab}=D_{a}\delta_{ab}$ and $P_{ab}=\varphi_{a}\psi_{b}$ be a $1$-dimensional projector. Then
(92) $\det(D-z-P)=\prod_{a}(D_{a}-z)-\sum_{\beta}\varphi_{b}\psi_{b}\prod_{\begin{subarray}{c}a\neq b\end{subarray}}(D_{a}-z)$
Proof.
We have
(93) $\det(D-z-P)$ $=\exp\left(\mathrm{Tr}\ln(D-P-z)\right)$
(94) $=\exp\left(\mathrm{Tr}\ln(D-z)+\operatorname{Tr}\ln\left(1-(D-z)^{-1}P\right)\right)$
(95) $=\prod_{\alpha}(D_{\alpha}-z)\exp\left(\sum_{k\geq 1}\frac{1}{k}\mathrm{Tr}\left((D-z)^{-1}P\right)^{k}\right)$
Now
(96) $\mathrm{Tr}\left((D-z)^{-1}P\right)^{k}$ $=\sum_{a_{1},\ldots,a_{k}}(D_{a_{1}}-z)^{-1}\cdots(D_{a_{n}}-z)^{-1}P_{a_{1}a_{2}}\cdots P_{a_{k}a_{1}}$
(97) $=\sum_{a_{1},\ldots,a_{k}}(D_{a_{1}}-z)^{-1}\cdots(D_{a_{k}}-z)^{-1}\prod_{i=1}^{k}\varphi_{a_{i}}\psi_{a_{i}}$
From here:
$\exp\left(\sum_{k=1}^{\infty}\frac{1}{k}\mathrm{Tr}\left((D-z)^{-1}P\right)^{k}\right)=1-\sum_{b=1}^{n}(D_{b}-z)^{-1}\varphi_{b}\psi_{b}$
Together with (93) this implies (92). ∎