### v2

# Low–dimensional tori in Calogero–Moser–Sutherland systems

Abstract.
The main result of this paper is an explicit description of the stratification of the phase space of Calogero-Moser-Sutherland integrable systems corresponding to Lie groups $SU(n)$. Each stratum is isomorphic to $\mathbb{T}^{s}\times\mathbb{R}^{s}$ where $s=0,1,\dots,n-1$. Tori in such strata are Liouville tori. The zero-dimensional stratum is the fixed point of the multi-time dynamics of the model.
1. Introduction
Recall that Calogero–Moser–Sutherland (CMS) system [1, 6, 8] is a model of $n$ one–dimensional particles on a circle. The phase space of this model is $T^{*}H_{\rm reg}$ where $H_{\rm reg}$ is the regular part of the Cartan subgroup of $SU(n)$, i.e. diagonal unitary matrices with distinct eigenvalues. The space $T^{*}H_{\rm reg}$ has natural coordinates $(p_{i},\gamma_{i})$, $p_{i}\in\mathbb{R}$, $\lvert\gamma_{i}\rvert=1$, $\sum_{i=1}^{n}p_{i}=0$, $\prod_{i=1}^{n}\gamma_{i}=1$, $\gamma_{i}\neq\gamma_{j}$.
The standard symplectic structure on $T^{*}H_{\rm reg}$ gives the following Poisson brackets:
(2) $\{p_{i},p_{j}\}=0,\quad\{p_{i},q_{j}\}=\delta_{ij},\quad\{q_{i},q_{j}\}=0.$
Here we used coordinates $q_{j},\gamma_{j}=e^{iq_{j}}$.
The Hamiltonian of the system is
(3) $H_{2}=\frac{1}{2}\sum_{i=1}^{n}p_{i}^{2}+\frac{1}{4}\sum_{i\neq j}^{n}\frac{c^{2}}{\sin^{2}\left(\tfrac{q_{i}-q_{j}}{2}\right)}.$
This is an integrable system with higher integrals
(4) $H_{k}=\frac{1}{k}\mathrm{Tr}(x^{k}),$
where the matrix $x$ as a function of $p$ and $q$ is given by formula (18). The matrix $x$ is known as the Lax matrix for the CMS system. Functions $\{H_{k}\}_{k=2}^{n}$ are independent [^1] They are independent in the usual sense: $dH_{2}\wedge\dots\wedge dH_{n}$ is not identically zero. and Poisson commute. This follows immediately from the Hamiltonian reduction construction [4]. Functions $H_{k}$ are invariant with respect to simultaneous permutations of $p$ and $q$. Thus, the effective phase space is a manifold $S(\mathcal{O})=T^{*}H_{reg}/S_{n}$ [^2] The notation $S(\mathcal{O})$ for the phase space is natural from the point of view of the Hamiltonian reduction, see section 2.1, where $S_{n}$ acts on $\mathfrak{h}^{*}\times H_{reg}$ diagonally.
Geometrically, the CMS system is a Lagrangian fibration $\pi:S(\mathcal{O})\to\mathfrak{h}/S_{n}$. Here $\mathfrak{h}\simeq\mathbb{R}^{n-1}$ is the Cartan subalgebra of $su(n)$ and $\pi([(x,g)])=[x]$, where $[(x,g)]$ and $[x]$ are conjugacy classes of $(x,g)$ and of $x$ respectively, see section 2.1 for details. One of the main results of this paper is an explicit description of $B(\mathcal{O})=Im(\pi)$ as a stratified space which is the shifted principal Weyl chamber for $su(n)$:
(5) $B(\mathcal{O})=\sqcup_{\{j\}}B(\mathcal{O})_{\{j\}}.$
Here $\{j\}=\{j_{0},j_{1},\dots,j_{s}\}$, where $j_{0}=1<j_{1}<\dots<j_{s}$, and $0\leq s\leq n-1$. The stratum $B(\mathcal{O})_{\{j\}}$ is isomorphic to $\mathbb{R}^{s}$. On this stratum $x_{i}-x_{i+1}=c$ if $i$ is not in the set $\{j\}$ and $x_{i}-x_{i+1}>c$ if $i\in\{j\}$. The maximum stratum $B(\mathcal{O})_{\{1,2,\dots,n\}}$ has dimension n1n-1 and is the interior of the shifted principal Weyl chamber, where $x_{i}-x_{i+1}>c$ for all $i=1,\dots,n-1$. Other strata are its walls.
As a consequence the phase space $S(\mathcal{O})$ gets a stratification by the preimages of $B(\mathcal{O})_{\{j\}}$.
(6) $S(\mathcal{O})=\sqcup_{\{j\}}S(\mathcal{O})_{\{j\}},$
where $S(\mathcal{O})_{\{j\}}=\pi^{-1}(B(\mathcal{O})_{\{j\}})$. In section 3.3 we prove that
(7) $S(\mathcal{O})_{\{j\}}\simeq\mathbb{T}^{s}\times\mathbb{R}^{s}.$
The maximal stratum $S(\mathcal{O})_{\{1,2,\dots,n\}}$ is subject to the Arnold–Liouville Theorem and its structure is well known, see for example [7]. The action-angle variables on this stratum are described in section 2.3. Action-angle variables on low–dimensional strata are described in section 3.3.
The minimal stratum $S(\mathcal{O})_{\{1\}}$ is the fixed point of the multi-time Hamiltonian flow generated by Poisson commuting Hamiltonians $H_{k}$ (see [5]).
$x_{2}-x_{3}$$x_{1}-x_{2}$cc22-dim11-dim11-dim0-dim
Figure 1. The shifted principal Weyl chamber for $G=SU(3)$ and dimensions of Liouville tori over the corresponding strata.
This paper is closely related to the work [5] on hybrid integrable systems and to the upcoming paper [2] on stratified super-integrable systems.
Similar stratification exists in many other models, for example, in the elliptic CMS system and integrable systems on moduli spaces of flat connections. We will address this in a separate publication. We also do not discuss here the algebro-geometric setting.
Acknowledgment
We are grateful to L. Feher for stimulating discussions and pointing out reference [3], to K. Jiang, Z. Chen, and H. Xiao for many discussions on stratified integrable systems.
The work of A. L. was supported by the Beijing Natural Science Foundation (IS24006) and Beijing Talent Program. The work of N. R. was supported by the Collaboration Grant ”Categorical Symmetries” from the Simons Foundation, by the Changjiang fund, and by the project 075-15-2024-631 funded by the Ministry of Science and Higher Education of the Russian Federation.
2. The Calogero–Moser–Sutherland system
2.1. The Hamiltonian reduction
Here we review some facts about the CMS system for $SU(n)$. From now on, $G$ is always $SU(n)$. This section is included for completeness of the exposition. It is a well-known $T^{*}G$ version of the Kazhdan-Kostant-Sternberg construction [4] of the Calogero-Moser system by the Hamiltonian reduction of $T^{*}\mathfrak{g}$.
Here and below $\mathfrak{g}=su(n)$. We treat elements of $su(n)$ as traceless $n\times n$ Hermitian matrices, and elements of $SU(n)$ as $n\times n$ unimodular unitary matrices.
The phase space for the CMS system for $G=SU(n)$ can be realized as the Hamiltonian reduction of $T^{*}G\simeq\mathfrak{g^{*}}\times G$ with respect to the Hamiltonian action of $G$ by conjugations. From now on, we identify $\mathfrak{g}\simeq\mathfrak{g}^{*}$ via the Killing form. For $su(n)$ this means that both $su(n)$ and $su(n)^{*}$ are represented by traceless Hermitian $n\times n$ matrices.
The Hamiltonian reduction for the adjoint action of $SU(n)$ on $T^{*}SU(n)$ via the coajoint orbit $\mathcal{O}\subset su(n)^{*}$ produces the symplectic leaf of the Poisson space $T^{*}SU(n)/SU(n)$ which is
(8) $\mathcal{S}(\mathcal{O})=\mu^{-1}(\mathcal{O})/G=\{(x,g)\in su(n)\times SU(n)\mid x-g^{-1}xg\in\mathcal{O}\}/G.$
Here $\mu:T^{*}G\to\mathfrak{g}^{*}$ is the moment map for the adjoint action, $\mu(x,g)=x-g^{-1}xg$.
When $\mathcal{O}$ is a coadjoint orbit of rank one, the corresponding Hamiltonian reduction is isomorphic to $T^{*}H_{\mathrm{reg}}/S_{n}$ as a symplectic manifold. Here $H_{\mathrm{reg}}=\{\mathrm{diag}(\gamma_{1},\ldots,\gamma_{n});\ \gamma_{i}\neq\gamma_{j}\}$ [^3] Clearly $T^{*}H_{\mathrm{reg}}/S_{n}$ is a smooth symplectic manifold..
A rank $1$ orbit in $su(n)$ consists of traceless Hermitian matrices of rank $1$. They can be parametrized as
(9) $\mu_{ij}=\overline{\varphi_{i}}\varphi_{j}-c\delta_{ij},\quad c=\frac{1}{n}\sum_{i=1}^{n}|\varphi_{i}|^{2}$
modulo transformations $\varphi_{j}\mapsto e^{i\theta}\varphi_{j}$. It is clear that the second Casimir function  [^4] The values of higher Casimir functions are the rank $1$ orbit are determinated by $c$ and (10) $\mathrm{Tr}(\mu^{k})=c^{k}\,(n-1)\left((n-1)^{k-1}+(-1)^{k}\right).$ is
(11) $\mathrm{tr}(\mu^{2})=c^{2}n(n-1).$
The moment map for the adjoint action of $G$ on brings $(x,g)\in T^{*}G$ to
(12) $\mu=x-g^{-1}xg.$
Functions
(13) $H_{k}(x,g)=\frac{1}{k}\mathrm{Tr}\big(x^{k}\big),\ \ k=2,\dots,n$
on $T^{*}G$ Poisson commute. This follows immediately from the symplectic structure on $T^{*}G$. Because they are $G$-invariant, they are functions on $T^{*}G/G$. As a consequence, their restriction to $S(\mathcal{O})=\mu^{-1}(\mathcal{O})/G\subset T^{*}G/G$ also Poisson commute. Functions $\{H_{k}\}_{k=2}^{n}$ are independent on $S(\mathcal{O})$ and because $\mathrm{dim}\,S(\mathcal{O})=2n-2$ they define a Liouville integrable system.
2.2. The diagonalization of $g$ and natural coordinates
Let us recall how to obtain the Hamiltonians of the CMS system.
Choose a representative of the SU(n)SU(n)–orbit through $(x,g)$ in (8) such that $g$ is diagonal:
(14) $g=\mathrm{diag}(\gamma_{1},\ldots,\gamma_{n}).$
The equation (12) for diagonal $g$ becomes
(15) $x_{ij}(1-\gamma_{i}^{-1}\gamma_{j})=\overline{\varphi}_{i}\varphi_{j},\quad$ $i\neq j,$
(16) $0=|\varphi_{i}|^{2}-c,\quad\qquad\qquad$ $i=j.$
Thus $|\varphi_{i}|^{2}=c,\ c>0$.
Note that if $\gamma_{i}=\gamma_{j}$, the equation (15) implies $\overline{\varphi}_{i}\varphi_{j}=0$ but this contradicts to $\lvert\varphi_{i}\rvert^{2}=c$. Thus, $\gamma_{i}\neq\gamma_{j}$.
When $\gamma_{i}\neq\gamma_{j}$ the stabilizer of (14) is the Cartan subgroup of $SU(n)$, i.e. the subgroup of unitary diagonal unimodular matrices. By the action of this subgroup on $x$ we can bring $\varphi_{j}$ to the form $\varphi_{j}=\sqrt{c}\,e^{i\theta}$. This describes the phase space as the quotient of $T^{*}H_{reg}$ (with coordinates $x_{ii}$ on $\mathfrak{g}^{*}$, and $\gamma_{i}$ on $H_{reg}\subset G$) with respect to the action of $S_{n}$ by permutation of pairs $(x_{ii},\gamma_{i})$.
For $j\neq k$ we have
(17) $x_{jk}=\frac{c}{1-\gamma_{j}^{-1}\gamma_{k}}.$
Setting $p_{i}:=x_{ii}$ we have
(18) $x=\begin{pmatrix}p_{1}&\frac{c}{1-\gamma_{1}^{-1}\gamma_{2}}&\ldots&\frac{c}{1-\gamma_{1}^{-1}\gamma_{n}}\\ \frac{c}{1-\gamma_{2}^{-1}\gamma_{1}}&p_{2}&\ldots&\ldots\\ \ldots&\ldots&\ldots&\ldots\\ \frac{c}{1-\gamma_{n}^{-1}\gamma_{1}}&\ldots&\ldots&p_{n}\end{pmatrix}$.
The tracelessness of $x$ implies $\sum_{i=1}^{n}p_{i}=0$, and the unimodularity of $\gamma$ implies $\gamma_{1}\dots\gamma_{n}=1$. For the first two Hamiltonians, we have:
(19) $H_{2}=\frac{1}{2}\mathrm{Tr}\big(x^{2}\big)=\frac{1}{2}\sum_{i=1}^{n}p_{i}^{2}+\frac{1}{2}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}\frac{c^{2}}{(1-\gamma_{j}^{-1}\gamma_{i})(1-\gamma_{i}^{-1}\gamma_{j})}=\frac{1}{2}\sum_{i=1}^{n}p_{i}^{2}+\frac{1}{4}\sum_{\begin{subarray}{c}i<j\end{subarray}}^{n}\frac{c^{2}}{\sin^{2}(\tfrac{q_{i}-q_{j}}{2})},$
(20) $H_{3}=\frac{1}{3}\mathrm{Tr}\big(x^{3}\big)=\frac{1}{3}\sum_{i=1}^{n}p_{i}^{3}+\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}\frac{c^{2}\,p_{i}}{(1-\gamma_{j}^{-1}\gamma_{i})(1-\gamma_{i}^{-1}\gamma_{j})}=\frac{1}{3}\sum_{i=1}^{n}p_{i}^{3}+\frac{1}{4}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}\frac{c^{2}\,p_{i}}{{\sin^{2}(\tfrac{q_{i}-q_{j}}{2})}}.$
Here $\gamma_{j}=\exp(iq_{j})$. The Hamiltonian reduction and the Poisson structure on $T^{*}G$ imply that $G$-invariant functions on $T^{*}G\simeq\mathfrak{g}^{*}\times G$ that come from $G$-invariant functions on $\mathfrak{g}^{*}$ form a Poisson commutative subalgebra.
2.3. The CMS system as a Lagrangian fibration
Geometrically, any integrable system is a projection from a 2n2n–dimensional symplectic manifold to an nn–dimensional space, such that the generic fiber is Lagrangian. For the CMS system this projection is
(21) $\pi:\mathcal{S}(\mathcal{O})\to{su(n)^{*}}/SU(n)\simeq\mathfrak{h}^{*}/W.$
Here $\mathcal{S}(\mathcal{O})=\mu^{-1}(\mathcal{O})/G$ is the phase space obtained by the Hamiltonian reduction, and $\mathfrak{h}\subset{su(n)}$ is the Cartan subgroup (real traceless diagonal matrices), and the Weyl group $W$ is the symmetric group $S_{n}$. From the results of the previous section, the CMS Lagrangian projection is equivalent to
(22) $\pi\colon T^{*}H_{\mathrm{reg}}\to\mathfrak{h}^{*}/W,\quad(p,\gamma)\mapsto[x]\in\mathfrak{g}^{*}/G\simeq\mathfrak{h}^{*}/W.$
Here $x$, as the function of $p$ and $\gamma$, is given by (18). For $su(n)$ we can identify $[x]$ with an ordered collection $(x_{1}\geq x_{2}\geq\dots\geq x_{n})$ of eigenvalues of $x$:
(23) $\mathfrak{h}^{*}/W\simeq\left\{x\in\mathbb{R}^{n}\ \middle|\ x_{1}+\dots+x_{n}=0,\ x_{1}\geq\dots\geq x_{n}\right\}\subset\mathbb{R}^{n-1}.$
This is the principal Weyl chamber for $su(n)$. Now let us describe the image of this projection as a shifted principal Weyl chamber in $\mathbb{R}^{n-1}$ and the stratification of $T^{*}H_{\mathrm{reg}}$ by the primages of walls of the shifted principal Weyl chamber.
3. The base of the Lagrangian projection for the CMS model and the stratification of $S(\mathcal{O})$
3.1. Solutions to the moment map equation
Here we will describe the image $B(\mathcal{O})\subset\mathfrak{h}^{*}/W$ of the Lagrangian projection (22). We show that it is a convex set which can be regarded as a stratified space, which is a shifted principal Weyl chamber. We also describe the induced stratification of the phase space $S(\mathcal{O})$.
Any $x\in su(n)^{*}$ is diagonalizable by the conjugation action of $SU(n)$. Let us choose a representative $(x,g)\in\mathfrak{g}^{*}\times G$ of the point $G(x,g)\in\mu^{-1}(\mathcal{O})/G\subset T^{*}G/G$ in the phase space where $x$ is diagonal:
(24) $x=\mathrm{diag}\left(x_{1},\ldots,x_{n}\right)\in\mathfrak{h}^{*}\subset su(n)^{*},\qquad x_{1}+x_{2}+\ldots+x_{n}=0,\qquad x_{1}\geq x_{2}\geq\dots\geq x_{n}.$
Theorem 1.
Solutions to the momentum map equation [^5] From now on, when it is not ambiguous, we will write $a$ instead of $a\cdot\text{Id}$ for the matrix which is a multiple of the identity matrix.
(25) $x-g^{-1}xg=\bar{\psi}\psi^{t}-c\ \$
can be described in the following way. Define $\lambda=g\bar{\psi}$, then the following holds:
1) The following conditions are equivalent
  • $x_{i}-x_{i+1}\geq c$ for all $i=1,\ldots,n-1$,
  • $|\psi_{i}|^{2}\geq 0$ for all $i=2,\ldots,n$,
  • $|\lambda_{i}|^{2}\geq 0$ for all $i=1,\ldots,n-1$.
Moreover, $|\psi_{1}|^{2}>0$ and $|\lambda_{n}|^{2}>0$.
2) For any $n>k\geq 1$, the following occurs only at the same time
  • $x_{k}-x_{k+1}=c$,
  • $|\psi_{k+1}|^{2}=0$,
  • $|\lambda_{k}|^{2}=0$.
3) If $x_{i}-x_{i+1}>c$ for $i=1,\ldots,n-1$, then
(26) $|\psi_{i}|^{2}=c\prod_{j=1,j\neq i}^{n}\frac{x_{i}-x_{j}+c}{x_{i}-x_{j}},\qquad\text{and }\qquad|\lambda_{i}|^{2}=c\prod_{j=1,j\neq i}^{n}\frac{x_{j}-x_{i}+c}{x_{j}-x_{i}}.$
4) For $x_{i}-x_{i+1}>c$ with $i=1,\ldots,n-1$ matrix $g$ has form
(27) $g_{ij}=\frac{\lambda_{i}\,\psi_{j}}{x_{j}-x_{i}+c}\ \ .$
Proof.
Let us start with part 1). The equation for the moment map can be written as
(28) $g^{-1}xg=x+c-\overline{\psi}\psi^{t}.$
Let us consider the equality of characteristic polynomials of both parts of this equation [^6] Compare with Lemma 3.4 from [3].,
(29) $\prod_{j=1}^{n}(x_{j}-z)=\mathrm{det}(x+c-\bar{\psi}\psi^{t}-z)=\prod_{j=1}^{n}(x_{j}+c-z)-\sum_{j=1}^{n}|\psi_{j}|^{2}\prod_{k\neq j}(x_{k}+c-z).$
Here we take into account that for a diagonal matrix AA
(30) $\mathrm{det}\left(A+uv^{t}\right)=\mathrm{det}\left(A\right)+v^{t}A^{\text{adj}}u,$
where $A^{\text{adj}}$ is the adjugate matrix to $A$, which defined by $A^{\text{adj}}A=\mathrm{det}A\cdot{\rm Id}$. This is a well-known fact which follows from lemma 1 in appendix A.
The identity (29) for $z=x_{i}+c$ is
(31) $\prod_{j=1}^{n}(x_{i}-x_{j}+c)=|\psi_{i}|^{2}\prod_{j\neq i}(x_{i}-x_{j}).$
which proves (26).
The equation (31) for $i=1$ implies that $|\psi_{1}|$ is strictly positive
(32) $|\psi_{1}|^{2}=c\prod_{j>1}\frac{x_{1}-x_{j}+c}{x_{1}-x_{j}}>0.$
In the equation (31) for $i=2$ the product over $j>2$ is strictly positive:
[…]
This implies that
(34) $|\psi_{2}|^{2}\geq 0$
if and only if $x_{1}\geq x_{2}+c$ and the equality $x_{1}=x_{2}+c$ is achieved if and only if $|\psi_{2}|^{2}=0$.
Continuing this reasoning, we get:
(35) $|\psi_{k}|^{2}\geq 0\ \$
if and only if $x_{k-1}\geq x_{k}+c$ and the equality $x_{k-1}=x_{k}+c$ is achieved only if $|\psi_{k}|^{2}=0$.
The equation for the moment map can be rewritten for $\lambda$ as
(36) $gxg^{-1}=x-c+\lambda\bar{\lambda}^{t}.$
The equality of characteristic polynomials gives
(37) $\prod_{j=1}^{n}(x_{j}-z)=\mathrm{det}(x-c+\lambda\bar{\lambda}^{t}-z)=\prod_{j=1}^{n}(x_{j}-c-z)+\sum_{j=1}^{n}|\lambda_{j}|^{2}\prod_{k\neq j}(x_{k}-c-z).$
Evaluating this idenity (37) at the point $z=x_{i}-c$ gives
(38) $\prod_{j=1}^{n}(x_{j}-x_{i}+c)=|\lambda_{i}|^{2}\prod_{j\neq i}(x_{j}-x_{i}).$
Using a similar argument as for equation (31), we get $|\lambda_{n}|^{2}>0$ and
(39) $|\lambda_{k-1}|^{2}\geq 0$
if and only if $x_{k-1}\geq x_{k}+c$, where equality $x_{k-1}=x_{k}+c$ is achieved only if $|\lambda_{k-1}|^{2}=0$. This proves part 1).
Part 2) directly follows from (31) and (38). We already proved 3).
Let us prove 4). After the multiplication of the momentum map (40) by $g$, the equation for the moment map becomes
(40) $gx+cg-xg=\lambda\psi^{t}\ \ .$
Matrix elements of this equation are a system of equations for $g_{ij}$:
(41) $(x_{j}-x_{i}+c)g_{ij}=\lambda_{i}\psi_{j},$
which implies (27). ∎
Corollary 1.
The image of the projection (22) is the subset $B(\mathcal{O})\subset\mathbb{R}^{n-1}/S_{n}$ which can be described as
(42) $B(\mathcal{O})=\{(x_{1},\dots,x_{n})\mid x_{i}-x_{i+1}\geq c,\ \sum_{i=1}^{n}x_{i}=0\}.$
This space is naturally stratified. Let $\{j\}=\{j_{0},j_{1},\dots,j_{s}\}$ be a collection where $j_{0}=1,1<j_{a}<j_{a+1}\leq n-1$. Then
(43) $B(\mathcal{O})=\sqcup_{\{j\}}B(\mathcal{O})_{\{j\}},$
where
[…]
It is easy to check that
(45) $\overline{B(\mathcal{O})_{\{j\}}}=\sqcup_{\{k\}\leq\{j\}}B(\mathcal{O})_{\{k\}}.$
Here $\{k\}\leq\{j\}$ means that $\{k\}$ is a subset of $\{j\}$. This stratification for $G=SU(3)$ is illustrated on Figure 1.
This stratification of $B(\mathcal{O})$ induces a stratification of $S(\mathcal{O})$ by the primages of $B$-strata, which we will describe below.
3.2. The maximal stratum
If $x_{i}>x_{i+1}+c$, the stabilizer of $x$ is the Cartan subgroup in $SU(n)$, i.e., unitary diagonal matrices. Acting by this stabilizer on $\psi$, we can bring it to the form $\psi_{i}=|\psi_{i}|$ for all $i$. After this, we have:
(46) $g=\alpha X,\qquad X_{ij}=\frac{|\lambda_{i}||\psi_{j}|}{x_{j}-x_{i}+c},\qquad\alpha_{ij}=\alpha_{i}\delta_{ij},\qquad|\alpha_{i}|=1,$
where $X$ is orthogonal matrix, i.e. $X^{t}=X^{-1}$. The Cauchy determinant formula implies that $\mathrm{det}\,X=1$, and therefore $\alpha_{1}\cdot\ldots\cdot\alpha_{n}=1$.
Remark 1.
We can multiply $g$ by a central element $\omega\in Z(SU(n))$:
(47) $\omega=\exp\left(\frac{2\pi i}{n}k\right){\rm Id},\qquad\text{ where }k=0,\ldots,n-1.$
If $\gamma_{j}$ are eigenvalues of the matrix $g$, such multiplication changes $\gamma_{j}\mapsto\omega\gamma_{j}$. Eigenvalues of $g$ are exponents of coordinates of the CMS model $\gamma_{j}=\exp(iq_{j})$. Thus the multiplication of $g$ by $\omega$ shifts coordinates as
(48) $q_{j}\mapsto q_{j}+\frac{2\pi}{n}k\ (\mathrm{mod}\ 2\pi)\ .$
Thus, we described the maximal stratum $S(\mathcal{O})_{\{1,2,\dots,n\}}\subset S(\mathcal{O})$. It has a natural coordinate system, $\{x_{i}\}$ that lie inside of the shifted principal Weyl chamber $x_{i}>x_{i+1}+c,\ \sum_{i=1}^{n}x_{i}=0$ and $\{\alpha_{i}\}$, $|\alpha_{i}|=1$ and $\prod_{i=1}^{n}\alpha_{i}=1$.
3.3. Low–dimensional strata $S(\mathcal{O})_{\{j\}}$
To describe all strata, we need the following technical fact.
Proposition 1.
If $\psi_{k}=0$ for some $k=2,\dots,n$, then the only nonzero matrix element of $g$ in the (k1)(k-1)-th row and the $k$-th column is $g_{k-1,k}$ and $|g_{k-1,k}|=1$.
Proof.
The equation (41) for $j=k$ becomes
(49) $(x_{k}-x_{i}+c)g_{ik}=0.$
If $g_{ik}=0$ for all $i$, then $g$ not invertible. Theorem 1 implies that $\psi_{k}=0$ if and only if $x_{k}-x_{k-1}+c=0$. Thus, in the equation above $g_{k-1,k}$ can be non-zero. For $j\neq k-1$ we have $x_{j}-x_{k}+c\neq 0$, and therefore $g_{jk}=0$. This means that the $k$-th column has only one non-zero element $g_{k-1,k}$.
The unitarity condition $\left(g^{{\dagger}}g\right)_{ik}=\sum_{j=1}^{n}\bar{g}_{ji}g_{jk}=\bar{g}_{k-1,i}g_{k-1,k}=\delta_{ik}$ implies that in the (k1)(k-1)-th row there is only one non zero element $g_{k-1,k}$, and $|g_{k-1,k}|=1$. ∎
Fix numbers $\{j\}=\{j_{0}=1,j_{1},\dots,j_{s}\}$ such that $1<j_{1}<j_{2}<\dots<j_{s}\leq n$.
Definition 1.
The stratum $S(\mathcal{O})_{\{j\}}\subset S(\mathcal{O})$ is the subspace where $\psi_{k}\neq 0$ when $k\in\{j_{0},j_{1},\dots,j_{s}\}$ and $\psi_{k}=0$ otherwise. Equivalently, $x_{k}-x_{k+1}>c$ if $k\in\{j_{0},j_{1},\dots,j_{s}\}$ and $x_{k}-x_{k+1}=c$ otherwise.
It is clear that $S(\mathcal{O})_{\{j\}}=\pi^{-1}(B(\mathcal{O})_{\{j\}})$.
Define $\widehat{g}=P^{-1}g$, where $P$ is the signed matrix of cyclic permutation
(50) $P=\begin{pmatrix}0&-1&0&\ldots&0\\ 0&0&-1&\ldots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\ldots&-1\\ 1&0&0&\ldots&0\\ \end{pmatrix}$.
Proposition 1 implies
(51) $\widehat{g}_{ij}=\delta_{ij}\alpha_{i},\quad|\alpha_{i}|=1$
if neither $i$ nor $j$ are not in $\{j_{a}\}_{a=0}^{s}$.
The moment map equation (12) give the following identity
(52) $x-\widehat{g}^{-1}x^{P}\widehat{g}=\overline{\psi}\psi^{t}-c,$
where $x^{P}=P^{-1}xP$ is the diagonal matrix with entries $x_{j}^{P}=x_{j-1}$ and $x_{0}=x_{n}$. For matrix elements of $\widehat{g}$ we obtain equations
(53) $\left(x_{j_{b}}-x^{P}_{j_{a}}+c\right)\widehat{g}_{j_{a}j_{b}}=\sum_{d=0}^{s}\widehat{g}_{j_{a}j_{d}}\,\overline{\psi}_{j_{d}}\,\psi_{j_{b}}.$
Define
(54) $\widehat{\lambda}_{j_{a}}=\sum_{d=0}^{s}\widehat{g}_{j_{a}j_{d}}\,\overline{\psi}_{j_{d}}.$
Theorem 2.
The following holds:
  • $\widehat{\lambda}_{i}=0$, if $i\neq j_{a}$, $a=0,\dots,s$.
  • (55) $|\psi_{j_{a}}|^{2}=\frac{\prod_{d=0}^{s}(x_{j_{a}}-x^{P}_{j_{d}}+c)}{\prod^{s}_{d\neq a}(x_{j_{a}}-x_{j_{d}})}.$
  • (56) $|\widehat{\lambda}_{j_{a}}|^{2}=\frac{\prod_{d=0}^{s}(x_{j_{d}}-x^{P}_{j_{a}}+c)}{\prod^{s}_{d\neq a}(x^{P}_{j_{d}}-x^{P}_{j_{a}})}.$
  • By the action of the Cartan part of the stabilizer of $x$, we can bring $\widehat{g}$ to the form
    (57) $\widehat{g}_{j_{a}j_{b}}=\alpha_{j_{a+1}}\frac{|\widehat{\lambda}_{j_{a}}|\,|\psi_{j_{b}}|}{x_{j_{b}}-x^{P}_{j_{a}}+c}.$
Proof.
Write the moment map equation as
(58) $\hat{g}^{-1}x^{P}\hat{g}=x-\overline{\psi}\psi^{t}+c.$
For the characteristic polynomial of the right side, using lemma 1 from appendix A we obtain
(59) $\prod_{d=0}^{s}(x_{j_{d}}^{P}-z)=\prod_{d=0}^{s}(x_{j_{d}}+c-z)-\sum_{b=0}^{s}|\psi_{j_{b}}|^{2}\prod_{\begin{subarray}{c}d=0\\ d\neq b\end{subarray}}^{s}(x_{j_{d}}+c-z).$
Evaluation of this identity at $z=x_{j_{a}}+c$ gives
(60) $\prod_{d=0}^{s}(x_{j_{d}}^{P}-x_{j_{a}}-c)=-|\psi_{j_{a}}|^{2}\prod_{\begin{subarray}{c}d=0\\ d\neq a\end{subarray}}^{s}(x_{j_{d}}-x_{j_{a}})$
which proves the formula (55) for $|\psi_{j_{a}}|^{2}$.
Similarly, writing the moment map equation as
(61) $\hat{g}\,x\,\hat{g}^{-1}=x^{P}+\widehat{\lambda}\,\widehat{\lambda}^{*}-c$
and using lemma 1 gives the proof of the formula (56) for $|\widehat{\lambda}_{j_{a}}|^{2}$.
Finally, the moment map equations (53) give
(62) $\widehat{g}_{j_{a}j_{b}}=\frac{\widehat{\lambda}_{j_{a}}\,\psi_{j_{b}}}{x_{j_{b}}-x_{j_{a}}^{P}+c}.$
By the action of the stabilizer of $x$, which is the Cartan subgroup $SU(n)$, we can make $\psi_{j_{b}}$ real. This gives formula (57). ∎
Note that to make $\psi_{j_{a}}$ real we need only a subgroup of the stabilizer of $x$.
Corollary 2.
We can choose a representative $(x,g)$ of the conjugacy class $(x,g)]\in\mu^{-1}(\mathcal{O})/G$, where $x$ is diagonal as above and $g$ is the product
(63) $g=\alpha\,P\,Y.$
Here $P$ is the cyclic permutation matrix (50), $Y$ is
$Y_{ij}$ $=\delta_{ij},\quad i,j\notin\{j_{0},\ldots,j_{s}\},$
(64) $Y_{i,j_{a}}=Y_{j_{a},i}$ $=0,\quad\,\ i\notin\{j_{0},\ldots,j_{s}\},$
(65) $Y_{j_{a}j_{b}}$ $=(-1)^{1-\delta_{a,0}}\frac{|\widehat{\lambda}_{j_{a}}||\psi_{j_{b}}|}{x_{j_{b}}-x_{j_{a}}^{P}+c},$
and $\alpha$ is the following diagonal matrix
(66) $\alpha_{ij}=\alpha_{i}\delta_{ij},\qquad|\alpha_{i}|=1,\qquad\text{and}\quad\alpha_{i}=1\ \text{ for }\ i\neq j_{a}.$
Indeed, let $\beta$ be an element of the stabilizer of $x$, i.e., a diagonal unitary matrix. We already used transformations
(67) $g\mapsto\beta g\beta^{-1}$
with $\beta_{i}=1$, when $i\neq j_{a}$, to make $\psi_{j_{a}}$ real. Let $\beta^{\prime}\in H$ be such that $\beta^{\prime}_{i}=1$, when $i=j_{a}$. We have
(68) $\beta^{\prime}g\beta^{\prime-1}=\beta^{\prime}\alpha PY\beta^{\prime-1}=\beta^{\prime}\alpha\beta^{\prime-1}_{P}PY.$
We can use these transformations to bring $\alpha$ to the form $\alpha_{i}=1$, $i\neq j_{a}$. This proves corollary 2.
One can show that
(69) $\mathrm{det}\,Y=1$
and that the real-valued matrix $Y$ is orthogonal
(70) $YY^{t}={\rm Id}.$
Note that because $\det(P)=1$, we have $\det(\alpha)=1$.
Corollary 3.
$S(\mathcal{O})_{\{j\}}\simeq\mathbb{T}^{s}\times\mathbb{R}^{s}$ with $x_{j_{a}}$ being coordinates on $\mathbb{R}^{s}$ and $\alpha_{j_{a}}$ being coordinates on $\mathbb{T}^{s}$.
3.4. The dynamical stratification of the phase space
The results of the previous sections imply that we have a stratification of $S(\mathcal{O})$ given by the preimages of walls of the shifted Weyl chamber:
[…]
The first part in this disjoint union is the fixed point $S(\mathcal{O})_{\{1\}}$ corresponding to $s=0$. Each stratum $S(\mathcal{O})_{\{j_{0},\dots,j_{s}\}}$ has dimension $2s$. The maximal stratum $S(\mathcal{O})_{\{1,2,\dots,n\}}$ has maximal dimension 2n22n-2.
This is a stratification where strata are parametrized by the partially ordered set $A=\{1\}\cup\dots\cup\{1<j_{1}<\dots<j_{s}\}$ with the partial order $\{k\}\leq\{j\}$ if $\{k\}$ is a subset of $\{j\}$. , see corollary 3, and
(72) $\overline{S(\mathcal{O})_{\{j\}}}=\sqcup_{\{k\}\leq\{j\}}S(\mathcal{O})_{\{k\}}.$
This property can be checked by taking the appropriate limit. For example, if $x_{i-1}\to x_{i}+c$ for $i\neq j_{a}$, $a=0,\dots,s$
(73) $|\psi_{i}|^{2}$ $\to 0,\qquad\qquad\text{for }i\neq j_{a},$
(74) $|\psi_{i}|^{2}$ $\to|\psi_{j_{a}}|^{2},\qquad\text{ for }i=j_{a}.$
Here on the left hand side $|\psi_{i}|^{2}$ is taken from (26), on the right hand side $|\psi_{j_{a}}|^{2}$ is taken from (55), and
(75) $|\lambda_{i}|^{2}$ $\to 0,\qquad\qquad\text{for }i\neq j_{a}-1,$
(76) $|\lambda_{i}|^{2}$ $\to|\widehat{\lambda}_{j_{a}}|^{2},\qquad\text{ for }i=j_{a}-1\ \mathrm{mod}\ n.$
3.4.1. Natural coordinates on $S(\mathcal{O})_{\{j\}}$
On the the maximum stratum $S(\mathcal{O})_{\{1,2,\dots,n\}}$ define variable $y_{i}=x_{i-1}-x_{i}$ for $j=2,\ldots,n$, such that $y_{i}\geq c$. Then
(77) $x_{i}=\frac{1}{n}\sum_{j=2}^{n}(n+1-j)y_{j}-\sum_{j=2}^{i}y_{j}.$
Define also $\theta_{j}=\tfrac{1}{i}\sum_{k=1}^{j-1}\ln\alpha_{k}$ for $j=2,\ldots,n$. Then,
(78) $\alpha_{k}=e^{i(\theta_{k+1}-\theta_{k})},\qquad k=1,\ldots n.$
Here we assume that $\theta_{1}=\theta_{n+1}=0$.
This gives a convenient set of coordinates on the maximal stratum: real coordinates $(y_{2},\ldots,y_{n})$ with constraints $y_{i}>c$ and coordinates $(\theta_{2},\ldots,\theta_{n})$ on unit tori. In section 4 we will show that $(y,\theta)$ are the action-angle variables.
On $S(\mathcal{O})_{\{j\}}$ define variables $y_{j_{a}}=x_{j_{a}}-x_{j_{a}-1}$ then
(79) $x_{i}=\frac{n-1}{2}c-(i-1)c+\sum_{a=1}^{s}\left(\frac{n+1-j_{a}}{n}-\delta_{i\geq j_{a}}\right)(y_{j_{a}}-c),$
where $\delta_{i\geq j}$ is a step function, $\delta_{i\geq j}=0$ if $i<j$ and $\delta_{i\geq j}=1$ if $i\geq j$.
Set $\alpha_{j_{a}}=e^{i\eta_{j_{a}}}$ and define $\theta_{j_{a}}=\sum_{b=a}^{s}\eta_{j_{b}}$. Thus, we described coordinates $(y_{j_{a}},\theta_{j_{a}})$ with $y_{j_{a}}>c$ on $S(\mathcal{O})_{\{j\}}$. We will show that these are action-angle variables on this stratum.
Because $\{y\}$ are coordinates on the base $B(\mathcal{O})_{\{j\}}$ of the Lagrangian fibration (22) of $S(\mathcal{O})_{\{j\}}$, they can be taken as action variables. In the next section we will show that $\{\theta\}$ are the corresponding angle variables.
3.4.2. Two examples of low–dimensional strata
Example 1.
The fixed point $S(\mathcal{O})_{\{1\}}$. In this case all $\psi_{i}=0$ for $i\geq 2$, i.e., $x_{i-1}-x_{i}=c$ for $i=2,\ldots,n$, and we have
  • $|\psi_{1}|^{2}=|\hat{\lambda}_{1}|^{2}=nc$.
  • $x_{i}=\frac{n-1}{2}c-(i-1)c$ for $1=2,\ldots,n$.
  • $g=P.$
In terms of original CMS coordinates we have $q_{j}=\frac{2\pi j}{n}$. The center of $SU(n)$, $g\mapsto\omega^{k}g$, $\omega^{n}=1$, acts by the cyclic shift $q_{j}\mapsto q_{j}+\frac{2\pi k}{n}({\rm mod}\ 2\pi)$.
Example 2.
Two–dimensional strata $S(\mathcal{O})_{\{1,j\}}$. Fix $2\leq j\leq n$ and assume that $\psi_{i}=0$ only when $i\neq 1,j$. The following is a consequence of theorem 2 and corollary 2
  • $|\psi_{1}|^{2}+|\psi_{j}|^{2}=|\hat{\lambda}_{1}|^{2}+|\hat{\lambda}_{j}|^{2}=nc$.
  • $x_{i-1}-x_{i}=c$ for $i=2,\ldots,n$ and $i\neq j$.
  • Define $y=x_{j-1}-x_{j}$, then $y>c$. Together with $\sum_{i=1}^{n}x_{i}=0$ this gives:
    (80) $x_{i}=\frac{n-1}{2}c-(i-1)c+\left(\frac{n+1-j}{n}-\delta_{i\geq j}\right)(y-c).$
  • $g=\alpha PY$, where $\alpha=\mathrm{diag}\left(\alpha_{1},1,\ldots,1,\alpha_{j}=\alpha_{1}^{-1},1,\ldots,1\right)$.
(81) $|\psi_{1}|^{2}$ $=(j-1)c\,\frac{y+(n-1)c}{y+(j-2)c},\qquad\ \quad|\psi_{j}|^{2}=(n-j+1)c\,\frac{y-c}{y+(j-2)c},$
$|\hat{\lambda}_{1}|^{2}$ $=(n-j+1)c\frac{y+(n-1)c}{y+(n-j)c},\qquad|\hat{\lambda}_{j}|^{2}=(j-1)c\ \frac{y-c}{y+(n-j)c},$
and
(82) $Y=\begin{pmatrix}r&0&\ldots&q&\ldots&0\\ 0&1&\ldots&0&\ldots&0\\ \vdots&\vdots&\ddots&\vdots&\ddots&\vdots\\ -q&0&\ldots&r&\ldots&0\\ \vdots&\vdots&\ddots&\vdots&\ddots&\vdots\\ 0&0&\ldots&0&\ldots&1\\ \end{pmatrix}\$,
where
(83) $r^{2}=\frac{(j-1)c\,(n-j+1)c}{(y+(j-2)c)(y+(n-j)c)},\qquad q^{2}=\frac{(y-c)(y+(n-1)c)}{(y+(j-2)c)(y+(n-j)c)}.$
Note that $r^{2}+q^{2}=1$, which implies that $\mathrm{det}(\,Y)=1$, therefore $\mathrm{det}\,\alpha=\alpha_{1}\,\alpha_{j}=1$.
4. Action-angle variables
Here we show that natural coordinates $(y,\theta)$ are Darboux action-angle variables on each $S(\mathcal{O})_{\{j\}}$.
The cotangent bundle $T^{*}G$ carries the natural symplectic form, which, after the trivialization of $T^{*}G$ by right translations can be written in terms of $(x,g)\in su(n)^{*}\times SU(n)\simeq T^{*}SU(n)$ as
(84) $\Omega=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}g\,g^{-1}\right).$
Here $x$ is an $n\times n$ Hermitian traceless matrix an $g$ is and unitary unimodular $n\times n$ matrix.
This symplectic form is $G$-invariant with respect to the lift of the conjugation action to $T^{*}G$. It acts on $(x,g)$ as
(85) $h\colon\quad x\mapsto hxh^{-1},\quad g\mapsto hgh^{-1}.$
The restriction of the symplectic form (84) to $\mu^{-1}(\mathcal{O})$ defines a presymplectic form, which remains $G$-invariant, and becomes a symplectic form $\Omega_{\mathcal{O}}$ on $S(\mathcal{O})=\mu^{-1}(\mathcal{O})/G$. We can model $S(\mathcal{O})$ as the space of pairs $(x,g)$, up to the action $H\subset SU(n)$, where $x=(x_{1}\geq x_{2}\geq\ldots\geq x_{n})$, $\sum_{i=1}^{n}x_{i}=0$, is diagonalized. On the maximal stratum $g=\alpha X$ (46), on other strata $g=\alpha PY$, see remark 2.
Now, let us describe the symplectic form $\Omega_{\mathcal{O}}$ in terms of $(y,\theta)$ coordinates introduced in section 3.4.1.
Proposition 2.
On the maximal stratum $S(\mathcal{O})_{\{1,2,\dots,n\}}$ the symplectic form $\Omega_{\mathcal{O}}$ is
(86) $\Omega_{\mathcal{O}}=-\sum_{j=1}^{n}\mathrm{d}x_{j}\wedge\mathrm{d}\eta_{j},$
where $x_{i}$ are eigenvalues of $x$ and $\alpha_{j}=\mathrm{exp}(i\eta_{j})$.
Proof.
Evaluating the symplectic form $\Omega$ on the representative $(x,g=\alpha X)$ we obtain
(87) $\Omega_{\mathcal{O}}=\left.i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}g\,g^{-1}\right)\right|_{\mu^{-1}(\mathcal{O})}=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}(\alpha X)\,X^{-1}\alpha^{-1}\right)=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}X\,X^{-1}\right)+i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}\alpha\,\alpha^{-1}\right).$
The first term vanishes:
(88) $\mathrm{Tr}\left(x\,\mathrm{d}X\,X^{-1}\right)=\sum_{i,j=1}^{n}x_{i}\,\mathrm{d}X_{ij}\cdot(X^{t})_{ji}=\frac{1}{2}\sum_{i,j=1}^{n}x_{i}\,\mathrm{d}(X_{ij})^{2}=\frac{1}{2}\sum_{i=1}^{n}x_{i}\,\mathrm{d}\sum_{j=1}^{n}(X_{ij})^{2}=0.$
Here we took into account the orthogonality of $X$, which implies $\sum_{j=1}^{n}(X_{ij})^{2}=1$. Thus, on the maximal stratum
(89) $\Omega_{\mathcal{O}}=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}\alpha\,\alpha^{-1}\right)=i\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\alpha^{-1}_{i}\mathrm{d}\alpha_{i}=-\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}\eta_{i}.$
After changing variables from $(x,\eta)$ to $(y,\theta)$ we can write $\Omega_{\mathcal{O}}$ as
(90) $\Omega_{\mathcal{O}}=-\sum_{j=2}^{n}\mathrm{d}y_{j}\wedge\mathrm{d}\theta_{j},$
where $y_{j}=x_{j-1}-x_{j}$ and $\theta_{j}=\sum_{i=1}^{j-1}\eta_{i}$. Thus, on the generic stratum $(y_{j},\theta_{j})$ are Darboux action-angle coordinates.
Proposition 3.
On the 2s2s–dimensional stratum $S(\mathcal{O})_{\{j\}}$ we have
(91) $\Omega_{\mathcal{O}}=-\sum_{a=1}^{s}\mathrm{d}y_{j_{a}}\wedge\mathrm{d}\theta_{j_{a}},$
where $y_{j_{a}}=x_{j_{a}-1}-x_{j_{a}}$ and $\theta_{j_{a}}=\sum_{b=a}^{s}\eta_{j_{b}}$, with $\alpha_{j_{b}}=e^{i\eta_{j_{b}}}$.
Proof.
Writing $g=\alpha PY$ as in (63) for the restriction of $\Omega_{\mathcal{O}}$ we get
(92) $\Omega_{\mathcal{O}}=\left.i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}g\,g^{-1}\right)\right|_{\mu^{-1}(\mathcal{O})}=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}(\alpha PY)\,Y^{-1}P^{-1}\alpha^{-1}\right)=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}Y\,Y^{-1}\right)+i\,\mathrm{d}\mathrm{Tr}\left(x^{P}\,\mathrm{d}\alpha\,\alpha^{-1}\right),$
where $x^{P}=P^{-1}xP$.
Define $\alpha_{j_{a}}=e^{i\eta_{j_{a}}}$ for $a=0,\ldots,s$. Constrain $\prod_{a=0}^{s}\alpha_{j_{a}}=1$ implies
(93) $\eta_{j_{0}}\equiv\eta_{1}=-\sum_{a=1}^{s}\eta_{j_{a}}.$
Expression $\mathrm{Tr}\left(x^{P}\,\mathrm{d}Y\,Y^{-1}\right)=0$ for the same reasons as in equation (88) and $\Omega_{\mathcal{O}}$ becomes
(94) $\Omega_{\mathcal{O}}=i\sum_{a=0}^{s}\mathrm{d}x^{P}_{j_{a}}\wedge\alpha_{j_{a}}^{-1}\mathrm{d}\alpha_{j_{a}}=$- $\sum_{a=1}^{s}\frac{1-j_{a}}{n}\mathrm{d}y_{j_{a}}\wedge\mathrm{d}\eta_{1}-\sum_{a=1}^{s}\sum_{b=1}^{s}\left(\frac{n+1-j_{b}}{n}-\delta_{j_{b}-1\geq j_{a}}\right)\mathrm{d}y_{j_{b}}\wedge\mathrm{d}\eta_{j_{a}}=$
$-\sum_{a=1}^{s}\sum_{b=1}^{s}\left(1-\delta_{j_{b}-1\geq j_{a}}\right)\mathrm{d}y_{j_{b}}\wedge\mathrm{d}\eta_{j_{a}}=-\sum_{b=1}^{s}\mathrm{d}y_{j_{b}}\wedge\sum_{a=b}^{s}\mathrm{d}\eta_{j_{a}}.$
Here we used (93). In term of $\theta_{j_{a}}=\sum_{a=b}^{s}\eta_{j_{b}}$ this gives (91). ∎
Thus, the coordinates $(y_{j_{a}},\theta_{j_{a}})$ are Darboux coordinates on $S(\mathcal{O})_{\{j\}}$. The coordinates $y_{j_{a}}$ are also coordinates on the stratum $B(\mathcal{O})_{\{j\}}$ of $B(\mathcal{O})$. Complementary coordinates $\theta_{j_{a}}$ are coordinates on the tori $\pi^{-1}(y)$ and are angle coordinates. Now let us show that the multi-time dynamics is linear in $\theta$.
5. The multi-time dynamics of Calogero–Moser–Sutherland system
5.1. The multi-time dynamics on $T^{*}G$
Let $F$ be a polynomial $G$-invariant function on $\mathfrak{g}^{*}$ and $v_{F}$ be the Hamiltonian vector field generated by the pull-back of $F$ to $T^{*}G\simeq\mathfrak{g}\times G$ with respect to the projection to the first factor. For $G=SU(n)$
(95) $v_{F}(x,g)=-i\sum_{k,j+1}^{n}(\nabla Fg)_{kj}\frac{\partial}{\partial g_{kj}}.$
Here $\nabla F(x)$ is the gradient of $F$. It is an $su(n)$-valued function on $su(n)^{*}$, defined as $(y,\nabla F(x))=\frac{d}{dt}F(x+ty)|_{t=0}$, where $(y,x)={\rm tr}(xy)$. The flow line of this vector field, passing theorugh $(x,g)$ at $t=0$ is
(96) $(x(t),g(t))=(x,e^{t\nabla F(x)}g)\ \ .$
Any two such functions $F_{1},F_{2}$ Poisson commute with respect to the standard Poisson bracket on $T^{*}G$, $\{F_{1},F_{2}\}=0$. As a consequence, the corresponding vector field commute $v_{F_{1}},v_{F_{2}}]=0$ and thus, for any $k$ such functions $F_{1},\dots,F_{k}$ the multi-time flow is is defined with flow lines
(97) $(x(\mathbf{t}),g(\mathbf{t}))=(x,e^{i\sum_{j=1}^{k}t_{j}\nabla F_{j}(x)}g)\ \ .$
This dynamics preserves the pull-back to $T^{*}G$ of any $G$-invariant function $H$ on $\mathfrak{g}^{*}$ and the moment map $\mu(x({\bf t}),g({\bf t}))=\mu(x,g)=x-g^{-1}xg$.
5.2. The multi-time dynamics on $S(\mathcal{O})_{\{j\}}$
As we proved in Corollary 2 on the stratum $S(\mathcal{O})_{\{j\}}$, where $\{j\}=\{j_{0}=1<j_{1}<\dots<j_{s}\leq n\}$, the matrix $g$ can be written as
(98) $g=\alpha PY.$
In coordinates $(y,\theta)$, described earlier, the restriction (the pull-back) of the symplectic form $\Omega$ to this stratum is
(99) $\Omega=-\sum_{a=1}^{s}\mathrm{d}y_{j_{a}}\wedge\mathrm{d}\theta_{j_{a}},$
which means that Poisson brackets between coordinate functions are:
(100) $\{y_{j_{a}},y_{j_{b}}\}=0,\qquad\{\theta_{j_{a}},\theta_{j_{b}}\}=0,\qquad\{\theta_{j_{a}},y_{j_{b}}\}=\delta_{ab}.$
For a function $F$ which is the pull-back to $T^{*}G$ of a $G$-invariant function on $\mathfrak{g}^{*}$, the Hamiltonian vector fields $v_{F}$ being restricted to $S(\mathcal{O})_{\{j\}}$ can be written as
(101) $v_{F}^{\{j\}}=\sum_{i=1}^{n}\sum_{a=1}^{s}\frac{\partial F}{\partial x_{i}}\frac{\partial x_{i}}{\partial y_{j_{a}}}\frac{\partial}{\partial\theta_{j_{a}}},\qquad\text{ for }k=2,\ldots,n.$
Using (77) we can express $x_{i}$ in terms of $y_{i}$. This gives
(102) $v_{F}^{\{j\}}=\sum_{a=2}^{s}u^{F,\{j\}}_{a}(y)\frac{\partial}{\partial\theta_{j_{a}}},$
where
(103) $u^{F,\{j\}}_{a}(y)=\frac{n+1-j_{a}}{n}\sum_{i=1}^{j_{a}-1}\frac{\partial F}{\partial x_{i}}(x(y))-\sum_{i=j_{a}}^{n}\frac{\partial F}{\partial x_{i}}(x(y)).$
Here the function $x(y)$ is given by (79).
As a consequence, the multi-time dynamics in $(y,\theta)$ coordinates is
(104) $y_{j_{a}}(\mathbf{t})=y_{j_{a}},\ \ \theta_{j_{a}}(\mathbf{t})=\theta_{j_{a}}-\sum_{k=1}^{s}u^{F_{k},\{j\}}_{a}(y)t_{k},$
where $\mathbf{t}=(t_{1},\dots,t_{s})$. The dynamics is linear in angle variables, as it should be.
Appendix A Lemma about the determinant of the sum of a matrix plus a rank $1$ matrix
Lemma 1.
Let $D$ be a diagonal matrix $D_{ab}=D_{a}\delta_{ab}$ and $P_{ab}=\varphi_{a}\psi_{b}$ be a 11–dimensional projector. Then
(105) $\det(D-z-P)=\prod_{a}(D_{a}-z)-\sum_{b}\varphi_{b}\psi_{b}\prod_{\begin{subarray}{c}a\neq b\end{subarray}}(D_{a}-z).$
Proof.
We have
(106) $\det(D-z-P)$ $=\exp\left(\mathrm{Tr}\ln(D-P-z)\right)$
(107) $=\exp\left(\mathrm{Tr}\ln(D-z)+\operatorname{Tr}\ln\left(1-(D-z)^{-1}P\right)\right)$
(108) $=\prod_{\alpha}(D_{\alpha}-z)\exp\left(\sum_{k\geq 1}\frac{1}{k}\mathrm{Tr}\left((D-z)^{-1}P\right)^{k}\right).$
Now
(109) $\mathrm{Tr}\left((D-z)^{-1}P\right)^{k}$ $=\sum_{a_{1},\ldots,a_{k}}(D_{a_{1}}-z)^{-1}\cdots(D_{a_{n}}-z)^{-1}P_{a_{1}a_{2}}\cdots P_{a_{k}a_{1}}$
(110) $=\sum_{a_{1},\ldots,a_{k}}(D_{a_{1}}-z)^{-1}\cdots(D_{a_{k}}-z)^{-1}\prod_{i=1}^{k}\varphi_{a_{i}}\psi_{a_{i}}.$
From here:
(111) $\exp\left(\sum_{k=1}^{\infty}\frac{1}{k}\mathrm{Tr}\left((D-z)^{-1}P\right)^{k}\right)=1-\sum_{b=1}^{n}(D_{b}-z)^{-1}\varphi_{b}\psi_{b}.$
Together with (106) this implies (105). ∎